### v1

# Reinforcement Fine-Tuning of Large Language Models

[Uncaptioned image]: A General-Purpose and Unified Framework for
Reinforcement Fine-Tuning of Large Language Models
Xuchen Pan, Yanxi Chen, Yushuo Chen, Yuchang Sun, Daoyuan Chen,
Wenhao Zhang, Yuexiang Xie, Yilun Huang, Yilei Zhang, Dawei Gao,
Yaliang Li, Bolin Ding, Jingren Zhou
Alibaba Group
11footnotetext: Equal contribution.22footnotetext: Corresponding author. {chenyanxi.cyx, panxuchen.pxc, yaliang.li, bolin.ding}@alibaba-inc.com
Trinity-RFT is a general-purpose, flexible and scalable framework designed for reinforcement fine-tuning (RFT) of large language models. It is built with a decoupled design, consisting of (1) an RFT-core that unifies and generalizes synchronous/asynchronous, on-policy/off-policy, and online/offline modes of RFT, (2) seamless integration for agent-environment interaction with high efficiency and robustness, and (3) systematic data pipelines optimized for RFT. Trinity-RFT can be easily adapted for diverse application scenarios, and serves as a unified platform for exploring advanced reinforcement learning paradigms. This technical report outlines the vision, features, design and implementations of Trinity-RFT, accompanied by extensive examples demonstrating the utility and user-friendliness of the proposed framework.
Note: Trinity-RFT is currently under active development. This technical report corresponds to commit id f17db3d (May 23, 2025) of the GitHub repository, and will be continuously updated as the codebase evolves. Comments, suggestions and contributions are welcome!
Figure 1: The design of Trinity-RFT.
1 Vision of Trinity-RFT
Reinforcement learning (RL) has achieved remarkable success in the development of large language models (LLMs). Examples include aligning LLMs with human preferences via reinforcement learning from human feedback (RLHF) [18], and training long-CoT reasoning models via RL with rule-based rewards [3, 28]. However, such approaches are limited in their abilities to handle dynamic, real-world, and continuous learning.
Trinity-RFT envisions a future where AI agents learn by interacting directly with environments, collecting delayed or complex reward signals, and continuously refining their behavior through RL based on the collected experiences [26]. For example, imagine an AI scientist who designs an experiment, executes it, waits for feedback (while working on other tasks concurrently), and iteratively updates itself based on true environmental rewards when the experiment is finally finished. Trinity-RFT offers a path into this future by providing various useful features.
2 Key Features
Trinity-RFT is a general-purpose, flexible, scalable and user-friendly RL framework that can be adapted for diverse experimental or real-world scenarios. The following features of Trinity-RFT make it a promising solution for realizing the aforementioned vision:
  • Unified RFT modes & algorithm support. Trinity-RFT unifies and generalizes existing RFT methodologies into a flexible and configurable framework, supporting synchronous/asynchronous, on-policy/off-policy, and online/offline training, as well as hybrid modes that seamlessly combine them into a single learning process (e.g., incorporating expert trajectories to accelerate an online RL process [15, 36]). This is made possible partly by our decoupled design that allows rollout and training to be executed separately and scaled up independently on different machines, which will soon be introduced in the next section.
  • Agent-environment interaction as a first-class citizen. Trinity-RFT allows delayed rewards in multi-step/time-lagged feedback loops, and handles long-tailed latencies and environment/agent failures gracefully, which ensures efficiency and robustness in real-world scenarios with complex agent-environment interaction.
  • Systematic data pipelines optimized for RFT. These include converting raw datasets to task sets for RFT, cleaning/filtering/prioritizing experiences stored in the buffer, incorporating priors and additional reward signals from various sources into the collected experiences, synthesizing data for augmenting tasks and experiences, offering user interfaces for human-in-the-loop, among others.
3 Design and Implementations
The overall design of Trinity-RFT exhibits a trinity consisting of (1) RFT-core, (2) agent-environment interaction, and (3) data pipelines, which are illustrated in Figure 1 and elaborated in this section.
3.1 RFT-Core
RFT-core is the component of Trinity-RFT, highlighted at the center of Figure 1, where the core RFT process happens. Its design also exhibits a trinity, consisting of the explorer, trainer, and buffer.
  • The explorer, powered by a rollout model, takes a task as input and solves it by executing a workflow that specifies the logic of agent-environment interaction, thereby collecting experiences (including rollout trajectories, rewards, and other useful information) to be stored in the buffer.
  • The buffer stores experiences that can be generated by the explorer or by other sources, such as human experts. It also assists with fetching training samples for the trainer, and can be integrated with advanced sampling strategies and post-processing operations.
  • The trainer, backed by a policy model, samples batches of experiences from the buffer and updates the policy model via RL algorithms.
Our implementations allow the explorer and trainer to be deployed on separate machines. This decoupled design of RFT-core enables Trinity-RFT to support diverse RFT modes with great flexibility, for example:
  • In a synchronous mode, the explorer and trainer get launched simultaneously, work in close coordination, and synchronize their model weights once every sync_interval training steps. Within each synchronization period, the explorer continuously generates sync_interval batches of rollout experiences and stores them in the buffer, which are then retrieved and utilized by the trainer for updating the policy model. If sync_interval $=1$, this is an on-policy RL process, whereas if sync_interval $>1$, it becomes off-policy (akin to the mode adopted in [28]) and can be accelerated by pipeline parallelism between the explorer and trainer. This mode can be activated by setting the configuration parameter mode to both.
  • In a fully asynchronous mode, the explorer and trainer are mostly decoupled and can scale up independently. The explorer continuously generates rollout experiences and stores them in the buffer, while the trainer continuously samples experiences from the buffer and uses them for training the policy model. External experiences, e.g., those generated by expert models or humans, can be continuously incorporated into the buffer as well. The explorer synchronizes its model weights with the trainer once in a while, keeping the distribution of rollout experiences up to date. This mode can be activated by setting mode to explore/train and launching the explorer and trainer separately on different machines.
  • Trinity-RFT supports a benchmark mode that allows the user to evaluate one or multiple checkpoints on arbitrary benchmarks, after the RFT training process has finished. To activate this mode, the user simply needs to set mode to bench and specify the paths for the evaluation datasets in the configurations. This mode can be particularly useful for experimental purposes; for example, the user might want to try out different RFT techniques or configurations quickly (with limited evaluation on hold-out data) during training, identify which RFT trials have achieved stable convergence and high rewards, and then conduct more thorough evaluations only for the checkpoints of these successful trials.
  • In certain scenarios, the user would like to train the policy model without further exploration, using experiences that have already been collected and stored in the buffer. This train-only mode can be activated by setting the mode parameter to train and launching the trainer alone. Offline methods like Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) [20] can be regarded as special cases of such scenarios, both of which are natively supported by Trinity-RFT. For another example, consider an online RFT process that expands over a long period of time, where the explorer alone is launched during the daytime for serving human users and collecting experiences, while the trainer alone is launched at night for updating the policy model (which will be thoroughly validated and evaluated before it can be actually deployed as the rollout model for the next day).
Another benefit brought by the decoupled design is that explorers and trainers can operate across separate devices (e.g., multiple explorers on edge devices, one trainer on a high-end GPU cluster) and scale up independently. Indeed, we anticipate that for successful RL in real-world scenarios, distributed collection of massive rollout experiences could be one promising solution to resolving data scarcity.
Remark 1.
We take a system-algorithm co-design perspective in the development of Trinity-RFT, and aim to unify and generalize diverse RFT methodologies in this framework, including synchronous/asynchronous, on-policy/off-policy, and online/offline modes. RFT-core provides the necessary infrastructure for achieving this goal. While the main text of this technical report focuses on the system perspective, interested readers may refer to Appendix A for further discussions on the algorithm perspective of off-policy/asynchronous RL.
3.2 Agent-Environment Interaction
Numerous challenges arise when one tries to build an RFT framework that can efficiently and robustly handle real-world interaction between the LLM-powered agent and the environment. These include long-tailed latencies, agent/environment failures, and lagged reward signals, among others.
Trinity-RFT regards agent-environment interaction as a first-class citizen and incorporates various solutions to tackle these challenges, for example:
  • Built upon vLLM [10], our implementations of the LLM inference engine in Trinity-RFT support asynchronous and concurrent generation of rollout trajectories for multiple tasks. This helps mitigate the straggler effect caused by the long-tailed latencies in rollout generation and agent-environment interaction, thereby accelerating the RFT process. Load balancing among multiple LLM inference engines within one RFT training course is also taken care of, and would be one direction for further optimizing the utilization of computational resources.
  • Trinity-RFT incorporates various timeout/retry/skip mechanisms for fault tolerance and robustness, which ensure that continuous rollout generation would not be interrupted or blocked by individual failures in certain rounds of agent-environment interaction. This is crucial for stable and efficient learning in real-world scenarios, e.g., when the agent interacts with a large number of MCP services [12] that differ vastly in quality and availability.
  • Trinity-RFT is built to provide native support for asynchronous RFT modes, which allow great flexibility in the paces of the explorer and trainer. This can boost the overall efficiency of the RFT process, compared to synchronous modes where the slower one among the explorer and trainer can block the progress of the other and cause waste of computational resources.
  • For lagged reward signals, the trinity design of RFT-core offers a natural solution. As soon as the rollout trajectory (without reward values) has been generated, it is saved into the experience buffer, but marked as “not ready for training”. The explorer is now free from this task and may continue to collect experiences for other tasks. When the reward signals from the environment finally arrive, they are written to the buffer, and the corresponding experience is now marked as “ready for training”.
To adapt Trinity-RFT to a new downstream scenario, the user mainly needs to define and register a customized workflow (by inheriting the base class Workflow or MultiTurnWorkflow) where the logic of agent-environment interaction is implemented. Detailed examples will be demonstrated in Sections 4.1 and 4.2.
3.3 Data Pipelines
The data pipelines in Trinity-RFT aim to address fundamental challenges in RFT scenarios, such as managing heterogeneous data dynamics across interaction workflows, enabling delayed reward integration, and facilitating continuous data curation. Our solutions center on three core aspects: end-to-end data transformation, active data-reward co-shaping, and human-in-the-loop curation, each corresponding to key requirements identified in our development of RFT-core (Section 3.1).
3.3.1 End-to-end Data Transformation
To support the diverse RFT modes (e.g., synchronous or asynchronous) in Trinity-RFT, we establish a service-oriented data pipeline architecture as illustrated in Figure 2. It decouples data pipeline logic from procedure control to enable flexible RL-oriented data transformations with two key modules:
  • The Formatter Module unifies disparate data sources into RFT-compatible formats, providing convenient conversion between raw inputs (e.g., meta-prompts, domain-specific corpora, and QA pairs with tagged rewards) and structured RFT representations (e.g., via RftDatasetModel). For efficient RFT workloads, we utilize PostgreSQL-based persistent storage supporting both ExperienceModel for prioritized trajectories and DPODataModel for preference pairs, which is particularly beneficial for asynchronous RFT mode and distributed environments. Full schema details appear in Appendix B.1.
  • The Controller Module manages the complete data pipeline lifecycle through distributed server initialization, declarative configuration, and automated dataset persistence. It implements dynamic control mechanisms for asynchronous scenarios and protection against resource exhaustion, with configurable termination conditions based on compute quota or data quantity. This modular design enables Trinity-RFT to handle data transformations flexibly while maintaining consistency across different RFT modes. More service utilities are detailed in Appendix B.2.
The Formatter-Controller duality mirrors the explorer-trainer decoupling in RFT-core, enabling parallel data ingestion and model updating. This design also allows Trinity-RFT to handle delayed rewards through version-controlled experience updates while maintaining low-latency sampling for the trainer.
Figure 2: An illustration of data flow in Trinity-RFT.
3.3.2 Active Data-Reward Co-Shaping
To enable efficient and effective learning from complex and delayed reward signals, Trinity-RFT introduces fruitful mechanisms that actively shape both data and reward throughout the RFT process.
  • Trinity-RFT provides the Processor Module to support convenient and extensive data processing capabilities tailored to RFT, through dedicated bridge classes that link Trinity-RFT to over 100 operators from the Data-Juicer system [2]. This module offers efficient integration with the experience buffer, and provides composable building blocks for experience cleaning (e.g., length filters, duplication removal), safety alignment (e.g., toxicity detection, ethics checks), and preference data synthesis (e.g., critique-conditioned augmentation). Users can specify processing chains using either natural language or explicit configurations. These are accessible to both RFT users familiar with Data-Juicer and those who are not, thanks to transparency of the underlying system details. By treating Data-Juicer as a modular data processing operator pool rather than a central dependency, Trinity-RFT provides RL-specific abstractions and coherence, while benefiting from well-established data tools.
  • RL-oriented data shaping. RL is particularly useful in scenarios where optimization goals cannot be clearly defined, leading users to express vague or high-level requirements. In response, Trinity-RFT automates the translation of these objectives into optimized data processing pipelines. It reduces problem complexity through a structured hierarchy: Quality $\rightarrow$ Difficulty $\rightarrow$ Diversity $\rightarrow$ Quantity, with auto-selection of appropriate operators from Data-Juicer. Moreover, Trinity-RFT provides many composable and extensible interfaces for data shaping, aiming to seamlessly inject user-defined inductive biases. For instance, DataCleaner (a sub-module of Processor) supports multi-way cleaning with statistical filtering (e.g., entropy-based diversity [11]) and rule-based validation (e.g., boxed answer verification in math dataset [7]), while DataSynthesizer enables both autonomous generation and hybrid augmentation with controlled feedback metrics [35].
  • Active data optimization. As agents interact with environments, experiences collected and stored in the buffer can become large-scale and highly diverse. Using the right experiences for RL training is crucial for a cost-effective and stable learning process. To this end, Trinity-RFT allows users to flexibly apply multi-dimensional utility scoring to prioritize samples (e.g., 0.4 * difficulty + 0.3 * diversity + 0.1 * quality - 0.2 * frequency). Moreover, we introduce the DataActiveIterator module to enable (1) version-controlled experience reuse through ExperienceModel; (2) cross-task data lineage tracking via rich relationships derived from sample_id and rollout-model_id; and (3) incorporating human annotations into the rewards. These features will be demonstrated through concrete examples in Section 4 and Appendix B.3.
3.3.3 Human-AI Collaboration
In scenarios where human feedback is irreplaceable, Trinity-RFT establishes a bi-directional human-AI collaboration loop that provides first-class support for human annotations, based on Label Studio [30] and Data-Juicer’s HumanOPs.
  • Multi-stage annotation. Trinity-RFT implements configurable procedures combining automatic pre-screening and human verification. Typical stages include preference annotation (comparative assessment of model responses), quality auditing (human verification of automated cleaning/synthesis results), and cold-start bootstrapping (initial dataset curation through expert demonstrations).
  • Native asynchronism support. As the collection of human feedback is generally slower than AI/model feedback, we provide dedicated capabilities to handle both synchronous and asynchronous feedback modes, with configurable timeout and polling parameters. The feedback collaboration is based on an event-driven design, with automatic task creation upon data state changes, configurable notifications via email/Slack/webhook, and an atomic transaction model for annotation batches.
  • Customization. Different applications may involve humans in heterogeneous ways. We thus prioritize flexibility in both the interaction-interface and service levels. Examples include rich built-in interfaces that can be extended in a visualized style with XML-like tags provided by Label Studio, fine-grained quality scoring for reward shaping, free-form feedback attachment for dataset shaping, among others. Moreover, for easy deployment, we provide local Label Studio instance management with automatic environment setup via Docker/pip; optimized SDK interactions with batch request coalescing; unified logging across annotation tools and ML services; and concurrent annotation campaigns through priority-based task routing, while maintaining full data lineage preserved via LineageTracker.
The decoupled design of Trinity-RFT, or the presence of an experience buffer in particular, enables human feedback to participate in RL loops without breaking the asynchronous execution model. For instance, human-verified samples can be prioritized for training while fresh experiences are being collected, which is a critical capability for real-world deployment scenarios with mixed feedback sources. Further details for human-AI collaboration in Trinity-RFT will be illustrated in Section 4.4.
3.4 Performance Optimizations and User-Friendliness
Trinity-RFT has made numerous efforts in pursuit of efficiency, for example:
  • For multi-turn conversations and ReAct-style workflows [37], Trinity-RFT supports concatenating multiple rounds of agent-environment interaction compactly into a single sequence, with proper masking that indicates which tokens need to be incorporated into the training objective of RL algorithms.
  • For the synchronous RFT mode, Trinity-RFT accelerates the RFT process via pipeline parallelism between the explorer and trainer.
  • Trinity-RFT utilizes Ray [14] for distributed runtime, and NCCL communication primitives [16] (whenever feasible) for model weight synchronization between the explorer and trainer.
In addition, Trinity-RFT has been designed with user-friendliness as a top priority, for example:
  • We include a monitor (built upon Wandb [32] and TensorBoard [29]) that allows the user to conveniently track the progress of an RFT process, both quantitatively (e.g., via learning curves for rewards and other metrics) and qualitatively (e.g., via concrete examples of rollout trajectories generated at different RL steps). See Figure 3 for an example snapshot of the monitor.
  • We also implement a configuration manager, as shown in Figure 4, that allows the user to create configuration files conveniently via a front-end interface.
  • We provide Trinity-Studio, an all-in-one unified UI that allows the user to configure and run data inspection, data processing, RFT learning process, etc., all by clicking the mouse and filling forms, without writing any code. More functions will be integrated in the future. An example for developing with Trinity-Studio will be introduced in Section 4.5.
Figure 3: A snapshot of the monitor implemented in Trinity-RFT.
(a) The “beginner” mode.
(b) The “expert” mode.
Figure 4: Snapshots of the configuration manager.
4 Examples and Applications
Through some concrete examples, we demonstrate how to use Trinity-RFT for different purposes, and also exemplifies some concepts introduced in previous sections. Additional step-by-step tutorials can be found on the documentation website [^1] https://modelscope.github.io/Trinity-RFT, or the /examples folder of the GitHub repository [^2] https://github.com/modelscope/Trinity-RFT.
4.1 Single-Turn Workflow
In a simple yet common scenario, a user of Trinity-RFT would like to train an LLM for completing single-turn tasks, where the LLM generates one response to each input query. For this purpose, the user mainly needs to (1) define and register a single-turn workflow class (by inheriting the base class Workflow) tailored to the targeted tasks, and (2) specify the dataset of tasks (for training and/or evaluation) and the initial LLM, both compatible with HuggingFace [9] and ModelScope [13] formats.
Listing 1 gives a minimal example for implementing a single-turn workflow. Suppose that each task is specified by a <question, answer> tuple. The run() method of ExampleWorkflow calls the LLM once to generate a response for the question, calculates its reward, and returns an Experience instance that consists of the response itself, the reward value, and the log-probabilities of next-token prediction by the rollout model (which is necessary for certain RL algorithms, such as PPO [22] and GRPO [23]). Some built-in workflows have been implemented in Trinity-RFT, e.g., the MathWorkflow class for math-related tasks.
Trinity-RFT also allows more flexible and advanced usage by changing just a few configuration parameters, e.g., running supervised fine-tuning (SFT) as a warmup stage before the RFT process, or choosing between the synchronous and asynchronous modes of RFT (as described in Section 3.1).
1@WORKFLOWS.register_module("example_workflow")
2class ExampleWorkflow(Workflow):
3
4 def __init__(self, model: ModelWrapper, task: Task, **kwargs):
5 super().__init__(model, task, **kwargs)
6 self.question = task.raw_task.get("question")
7 self.answer = task.raw_task.get("answer")
8
9 def calculate_reward(self, response: str, truth: str) -> float:
10 return 1.0 if response == truth else 0.0
11
12 def run(self) -> List[Experience]:
13 response = self.model.chat(
14 [
15 {
16 "role": "user",
[…]
18 }
19 ],
20 n=self.task.rollout_args.repeat_times,
21 temperature=self.task.rollout_args.temperature,
22 )
23 reward: float = self.calculate_reward(response.response_text, self.answer)
24 return [
25 Experience(
26 tokens=response.tokens,
27 prompt_length=response.prompt_length,
28 reward=reward,
29 logprobs=response.logprobs,
30 )
31 ]
Listing 1: A minimal example for implementing a customized workflow.
4.2 Multi-Turn Workflow
In more advanced cases, the user would like to train an LLM-powered agent that solves multi-turn tasks by repeatedly interacting with the environment. With Trinity-RFT, achieving this is mostly as simple as in the single-turn case, except that the user needs to define and register a multi-turn workflow class (by inheriting the base class MultiTurnWorkflow), where the logic of agent-environment interaction is implemented. Listing 2 provides one such example for ALFWorld [25]. For training efficiency, the process_messages_to_experience() method concatenates multiple rounds of agent-environment interactions compactly into an Experience instance consisting of a single token sequence with proper masking, which can readily be consumed by standard RL algorithms like PPO and GRPO.
1@WORKFLOWS.register_module("alfworld_workflow")
2class AlfworldWorkflow(MultiTurnWorkflow):
3 """A workflow for the ALFWorld task."""
4
5 # ...
6
7 def generate_env_inference_samples(self, env, rollout_num) -> List[Experience]:
8 print("Generating env inference samples...")
9 experience_list = []
10 for i in range(rollout_num):
11 observation, info = env.reset()
12 final_reward = -0.1
13 memory = []
14 memory.append({"role": "system", "content": AlfWORLD_SYSTEM_PROMPT})
15 for r in range(self.max_env_steps):
16 format_obs = format_observation(observation)
17 memory = memory + [{"role": "user", "content": format_obs}]
18 response_text = self.get_model_response_text(memory)
19 memory.append({"role": "assistant", "content": response_text})
20 action = parse_action(response_text)
21 observation, reward, done, info = env.step(action)
22 if done:
23 final_reward = reward
24 break
25 experience = self.process_messages_to_experience(
26 memory, final_reward, {"env_rounds": r, "env_done": 1 if done else 0}
27 )
28 experience_list.append(experience)
29 # Close the env to save CPU memory
30 env.close()
31 return experience_list
32
33
34 def run(self) -> List[Experience]:
35 # ...
36 game_file_path = self.task_desc
37 rollout_n = self.repeat_times
38 # ...
39 env = create_environment(game_file_path)
40 return self.generate_env_inference_samples(env, rollout_n)
Listing 2: An implementation of a multi-turn workflow for ALFWorld [25].
4.3 Dataset Curation
We present an example to showcase the end-to-end data transformation (Section 3.3.1) and data-reward co-shaping (Section 3.3.2) capabilities of Trinity-RFT in a mathematical reasoning task. As shown in Listings 3 and 4, the user configures dataset paths and field mappings through a centralized DataConfig, then specifies natural-language processing instructions via dj_process_desc. The system will automatically translate these requirements into executable pipelines through three phases: (1) Processor applies LLM-powered difficulty scoring using Qwen-Max; (2) Controller optimizes quality-diversity tradeoffs through adaptive filtering thresholds and thereby prioritizes high-value training samples; (3) Formatter converts raw arithmetic problems into an RL-ready task set with specified reward functions and workflows.
This procedure exemplifies our active data optimization capability (Section 3.3.2) by combining automated quality assessments with configurable curation policies. Users can extend this pattern to many other domains through customization and adaptation: for example, users can define their own difficulty metrics and calculation rules in the parameters of Data-Juicer’s OPs. Besides, users can flexibly utilize another Synthesizer module and extend new Controller policies to support seamless transitions between different domain-specific reward shaping and data optimization priors.
1# Core dataset configuration
2class DataConfig:
3 """Centralized dataset configuration management"""
4
5 def __init__(self):
6 self.data = {
7 "dataset_path": "/path/to/gsm8k",
8 "format_config": {
9 "prompt_key": "question",
10 "response_key": "answer"
11 },
12 "dj_process_desc": "Compute difficulty scores for math questions",
13 "agent_model_name": "qwen-max"
14 }
15
16 def get_metadata(self) -> dict:
17 """Retrieve configuration metadata"""
18 return self.data
Listing 3: Core dataset configuration, which includes processing instructions as metadata.
1# Typical procedure that transforms raw dataset through multi-stage pipeline
2my_dataset = RftDataset(data_config) # Raw dataset auto-loading
3
4my_dataset.format([
5 BoxedMathAnswerFormatter(config), # Domain-specific formatting
6 RLHFFormatter(config) # RLHF format conversion
7 ])
8
9my_dataset.to_taskset(
10 reward_fn=AccuracyReward, # Reward shaping specification
11 workflow=MathWorkflow # Execution environment binding
12 )
Listing 4: Dataset transformation with domain-specific formatting and RLHF conversion.
4.4 Human-AI Collaborative Annotation
This example demonstrates the human-in-the-loop capability in Trinity-RFT for preference modeling. As illustrated in Listing 5 and Figure 5, our framework integrates Label Studio’s annotation interface with asynchronous data pipelines through four coordinated stages: (1) task generation: auto-creating annotation batches from model rollouts; (2) interactive labeling: providing UI for side-by-side response comparison; (3) quality control: enforcing inter-annotator agreement thresholds; and (4) versioned storage: tracking preference lineage in pre-defined fields like those in DPODataModel.
This pipeline reflects Trinity-RFT’s bi-directional collaboration feature (Section 3.3.3), backed by timeout-aware task polling and support of atomic batch commit. It enables hybrid procedures where initial AI pre-screening can reduce human workload in production deployments. Annotation activities can scale across distributed teams through event-driven task routing. The system’s flexibility benefits rapid adaptation to diverse annotation protocols, allowing developers to implement custom labeling interfaces through XML-based templates or integrate third-party annotation services via unified SDK endpoints. This capability underpins advanced use cases such as safety red-teaming datasets and online instruction tuning scenarios where human judgment remains irreplaceable for quality-critical decisions, particularly in human-centric sociocultural contexts where data quality, difficulty, and reward signals are difficult to verify logically.
1# Human annotation configuration
2class HumanAnnotationConfig:
3 """Preference annotation pipeline configuration"""
4
5 def __init__(self):
6 self.process = [
7 {
8 "human_preference_annotation_mapper": {
9 "wait_for_annotations": True, # Block until annotations complete
10 "timeout": 3600, # Maximum wait time in seconds
11 "prompt_key": "prompt", # Source field for prompts
12 "answer1_key": "answer1", # First candidate response
13 "chosen_key": "chosen" # Selected response key
14 }
15 }
16 ]
17
18 def get_pipeline(self) -> List[Dict]:
19 """Get annotation processing pipeline"""
20 return self.process
Listing 5: Configuration for human preference annotation.
Figure 5: An interactive interface for human preference annotation.
4.5 Trinity-Studio
Trinity-Studio provides visual interaction for the core capabilities of Trinity-RFT, designed to bridge the gap between system complexity and user accessibility. As shown in Figure LABEL:fig:trinity-studio:dashboard, its three integrated modules, “Training Portal”, pgAdmin”, and “Label Studio”, form a cohesive interface that directly supports the active data-reward co-shaping and human-AI collaboration capabilities established in Section 3.3.
  • “Training Portal” (Figure LABEL:fig:trinity-studio:training) implements configuration-to-execution procedures through declarative YAML editing with live validation and auto completion. This design aligns with Trinity-RFT’s decoupled architecture, enabling users to optionally specify dataset paths, processing and training parameters, meanwhile ensuring schema-aware validation against pre-defined data schema constraints (Section B.1). The live validation mechanism prevents misconfigurations that could disrupt the explorer-trainer coordination described in Section 3.1. Furthermore, the integration of runtime metrics with tools like Wandb/TensorBoard directly helps the active data optimization feature by surfacing signals such as difficulty distribution drifts and diversity metrics mentioned in Section 4.3. This transparency ensures that users can monitor how data curation strategies impact RFT performance in real time.
  • “pgAdmin” (Figure LABEL:fig:trinity-studio:pgadmin) reflects Trinity-RFT’s end-to-end data transformation capabilities by providing a visual panel for PostgreSQL-based storage (Section 3.3.1). This design benefits the versioned data lineage requirements of RFT, particularly for scenarios involving asynchronous training (Section 3.3.2). With intuitive SQL query builders, users can easily adjust schema, audit training experiences and human annotation batches with fine-grained precision. This capability is valuable for rapid validation of active learning policies by cross-referencing training outcomes with metadata (e.g., difficulty scores and staleness in asynchronous mode).
  • “Label Studio” page (Figure LABEL:fig:trinity-studio:label-studio) operationalizes Trinity-RFT’s bi-directional human-AI collaboration capability (Section 3.3.3). Utilizing the provided task polling and atomic batch commit mechanisms, users can annotate the data or experiences directly, allowing an asynchronous way to involve human feedback and to dynamically influence data curation.
By unifying these capabilities in a single UI, Trinity-Studio reduces the cognitive load of managing complex RFT procedures. For example, a researcher tuning a math reasoning task could use the Training Portal to adjust difficulty scoring parameters, view the resulting distribution shifts in the pgAdmin module, and then validate human annotators’ preferences in the Label Studio page. This end-to-end visibility is critical for debugging and iterating on RFT strategies, particularly when balancing tradeoffs between quality, diversity, and reward shaping. Moreover, this visual operation complements the programmatic APIs of Trinity-RFT while maintaining full compatibility with CLI procedures. We implement Trinity-Studio with Singe-Spa framework [27]. The modular architecture enables custom view development through JavaScript plugins and flexible extensions for general-purpose usage.
Figure 6: Snapshots of Trinity-Studio.
5 Conclusion
This report has provided an overview of Trinity-RFT, a general-purpose, flexible, scalable and user-friendly framework for reinforcement fine-tuning of large language models. Trinity-RFT offers a path into “the era of experience” [26], by supporting applications in diverse scenarios with complex agent-environment interaction, and serving as a unified platform for exploring advanced RL paradigms.
Acknowledgements
Trinity-RFT is built upon many excellent open-source projects, including but not limited to: verl [24] and PyTorch’s FSDP [5] for LLM training; vLLM [10] for LLM inference; Data-Juicer [2] for data processing functionalities; AgentScope [6] for agentic workflow; and Ray [14] for distributed systems. We have also drawn inspiration from RL/RFT projects like OpenRLHF [8], TRL [31], ChatLearn [1], and TinyZero [19].
Appendix A Off-Policy Reinforcement Learning Algorithms
For off-policy learning, one might simply adopt offline methods like supervised fine-tuning and direct preference optimization (DPO) [20], or their iterative variants where (some of) the training samples are generated online by the policy model itself and filtered by their reward values, e.g., rejection-sampling fine-tuning and iterative DPO. There also exist more advanced RL algorithms with dedicated design for off-policy or asynchronous scenarios [15, 21, 28, 4, 17, 33, 34, 36].
In the rest of this section, we present some of our findings about off-policy RL (with a focus on the bandit setting) through the lens of online policy mirror descent (OPMD). At the end of this section, we arrive at the surprising conclusion that the standard policy gradient, weighted by a coefficient and using the group mean reward as the baseline, can be a feasible direction for updating the policy even in off-policy settings (while the standard theory of policy gradient only holds for on-policy settings). This has been validated empirically in our exploratory experiments during the development of Trinity-RFT.
A.1 OPMD: Kimi’s Version
Below is a recap of the OPMD variant proposed in the technical report of Kimi k1.5 [28].
Analysis.
For a specific task/query $x$ and a reference policy $\pi_{\textsf{ref}}$, consider the following objective for training policy $\pi_{\bm{\theta}}$ at a particular iteration of the RL process:
$\max_{\bm{\theta}}\quad J(\bm{\theta};x,\pi_{\textsf{ref}})\coloneqq{\mathbb{E}}_{y\sim\pi_{\bm{\theta}}(\cdot|x)}[r(x,y)]-\tau\cdot D_{\textsf{KL}}\big{(}\pi_{\bm{\theta}}(\cdot|x)\|\pi_{\textsf{ref}}(\cdot|x)\big{)}.$
Note that $\pi_{\textsf{ref}}$ can be changing during the RL process. In [28], $\pi_{\textsf{ref}}$ is set to $\pi_{\bm{\theta}_{t}}$ at the $t$-th iteration, i.e., when updating the policy from $\bm{\theta}_{t}$ to $\bm{\theta}_{t+1}$.
The optimal policy $\pi^{\star}$ for this objective satisfies the following: for any response $y$,
$\pi^{\star}(y|x)=\frac{\pi_{\textsf{ref}}(y|x)e^{r(x,y)/\tau}}{Z}\propto\pi_{\textsf{ref}}(y|x)e^{r(x,y)/\tau},$ (1)
$\text{where}\quad Z\coloneqq Z(x,\pi_{\textsf{ref}})=\int\pi_{\textsf{ref}}(y^{\prime}|x)e^{r(x,y^{\prime})/\tau}\mathop{}\!\mathrm{d}y^{\prime}={\mathbb{E}}_{y^{\prime}\sim\pi_{\textsf{ref}}(\cdot|x)}[e^{r(x,y^{\prime})/\tau}].$ (2)
Taking logarithm of both sides of Eq. (1), we see that the optimal policy $\pi^{\star}$ must satisfy the following consistency condition:
$r(x,y)-\tau\cdot\log Z-\tau\cdot\big{(}\log\pi^{\star}(y|x)-\log\pi_{\textsf{ref}}(y|x)\big{)}=0.$
Algorithm.
Based on the above analysis, [28] proposes the following OPMD variant. For a query $x$, first sample $K$ rollouts $y_{1},\dots,y_{K}\sim\pi_{\textsf{ref}}(\cdot|x)$ from the reference policy, then define a surrogate loss as follows:
$\widehat{J}(\bm{\theta};x,\pi_{\textsf{ref}})\coloneqq\sum_{i\in[K]}\Big{(}r(x,y_{i})-\tau\cdot\log\widehat{Z}-\tau\cdot\big{(}\log\pi_{\bm{\theta}}(y_{i}|x)-\log\pi_{\textsf{ref}}(y_{i}|x)\big{)}\Big{)}^{2},$
$\text{where}\quad\tau\cdot\log\widehat{Z}\coloneqq\tau\cdot\log\Big{(}\frac{1}{K}\sum_{i\in[K]}e^{{r(x,y_{i})}/{\tau}}\Big{)}.$
Although this is an off-policy method (since the rollout policy $\pi_{\textsf{ref}}$ is different from the policy $\pi_{\bm{\theta}}$ being updated), it is still limited because the rollouts have to be sampled from the particular policy $\pi_{\textsf{ref}}=\pi_{\bm{\theta}_{t}}$ for the $t$-th iteration of the RL process, as mentioned earlier. The reason for this limitation is the need of estimating $Z=Z(x,\pi_{\textsf{ref}})$ using samples from $\pi_{\textsf{ref}}(\cdot|x)$.
A.2 Pairwise OPMD
Analysis.
To eliminate the $Z$ term, we note that Eq. (1) is equivalent to the following:
$\forall\ y_{1}\text{ and }y_{2},\quad\frac{\pi^{\star}(y_{1}|x)}{\pi^{\star}(y_{2}|x)}=\frac{\pi_{\textsf{ref}}(y_{1}|x)}{\pi_{\textsf{ref}}(y_{2}|x)}e^{\big{(}r(x,y_{1})-r(x,y_{2})\big{)}/\tau}.$
Taking logarithm of both sides, we have
$\log\pi^{\star}(y_{1}|x)-\log\pi^{\star}(y_{2}|x)=\log\pi_{\textsf{ref}}(y_{1}|x)-\log\pi_{\textsf{ref}}(y_{2}|x)+\frac{r(x,y_{1})-r(x,y_{2})}{\tau},$
or equivalently,
$r(x,y_{1})-\tau\cdot\big{(}\log\pi^{\star}(y_{1}|x)-\log\pi_{\textsf{ref}}(y_{1}|x)\big{)}=r(x,y_{2})-\tau\cdot\big{(}\log\pi^{\star}(y_{2}|x)-\log\pi_{\textsf{ref}}(y_{2}|x)\big{)}.$
Note that this holds true for a pair of arbitrary responses $y_{1}$ and $y_{2}$.
Algorithm.
For a query $x$ and $K$ arbitrary responses $y_{1},\dots,y_{K}$, we define the following surrogate loss:
$\widehat{J}(\bm{\theta};x,\pi_{\textsf{ref}})\coloneqq\sum_{1\leq i<j\leq K}\big{(}a_{i}-a_{j}\big{)}^{2},$
$\text{where}\quad a_{i}\coloneqq r(x,y_{i})-\tau\cdot\big{(}\log\pi_{\bm{\theta}}(y_{i}|x)-\log\pi_{\textsf{ref}}(y_{i}|x)\big{)},\quad i\in[K].$
Here $\pi_{\textsf{ref}}$ can be any reference policy for KL regularization, regardless of how $y_{1},\dots,y_{K}$ were sampled. While this is a fully off-policy RL method, it has its own limitation: to run this algorithm, we should make sure that multiple (at least 2) rollouts for the same task are included within one micro-batch (whose size is typically much smaller that of a batch or mini-batch), which adds to infrastructure complexity.
Remark 2.
In the special case of $K=2$, the above method, termed as “pairwise OPMD”, turns out to be the same as “contrastive policy gradient” proposed in [4], albeit with a simpler and more intuitive derivation.
A.3 OPMD: an Embarrassingly Simple Variant
Analysis.
Consider the $t$-th iteration of the RL process, i.e., updating from $\bm{\theta}_{t}$ to $\bm{\theta}_{t+1}$, and use $\pi_{\textsf{ref}}=\pi_{\bm{\theta}_{t}}$ as the reference policy. For a specific task/query $x$, recall from Section A.1 the original objective:
$\max_{\bm{\theta}}\quad J(\bm{\theta};x,\pi_{\bm{\theta}_{t}})\coloneqq{\mathbb{E}}_{y\sim\pi_{\bm{\theta}}(\cdot|x)}[r(x,y)]-\tau\cdot D_{\textsf{KL}}\big{(}\pi_{\bm{\theta}}(\cdot|x)\|\pi_{\bm{\theta}_{t}}(\cdot|x)\big{)}.$
We leverage the analysis in Section A.2, and take a closer look at the following pairwise loss for $a_{i}$ and $a_{j}$, normalized by $1/(1+\tau)^{2}$ to make the loss scale invariant to the hyperparameter $\tau$:
$\frac{(a_{i}-a_{j})^{2}}{(1+\tau)^{2}}$ $=\frac{1}{(1+\tau)^{2}}\bigg{[}\Big{(}r(x,y_{i})-r(x,y_{j})\Big{)}$
$\quad\quad-\tau\cdot\Big{(}\big{(}\log\pi_{\bm{\theta}}(y_{i}|x)-\log\pi_{\bm{\theta}_{t}}(y_{i}|x)\big{)}-\big{(}\log\pi_{\bm{\theta}}(y_{j}|x)-\log\pi_{\bm{\theta}_{t}}(y_{j}|x)\big{)}\Big{)}\bigg{]}^{2}.$
The trick here is that, if we only intend to take one gradient step of this loss at $\bm{\theta}=\bm{\theta}_{t}$, then the value of $(\log\pi_{\bm{\theta}}(y_{i}|x)-\log\pi_{\bm{\theta}_{t}}(y_{i}|x))-(\log\pi_{\bm{\theta}}(y_{j}|x)-\log\pi_{\bm{\theta}_{t}}(y_{j}|x))$ is simply zero. As a result,
$\nabla_{\bm{\theta}}\frac{(a_{i}-a_{j})^{2}}{(1+\tau)^{2}}\Big{|}_{\bm{\theta}_{t}}=\frac{-2\tau}{(1+\tau)^{2}}\Big{(}r(x,y_{i})-r(x,y_{j})\Big{)}\Big{(}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x)|_{\bm{\theta}_{t}}-\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{j}|x)|_{\bm{\theta}_{t}}\Big{)},$
and thus
$\nabla_{\bm{\theta}}\sum_{1\leq i<j\leq K}\frac{(a_{i}-a_{j})^{2}}{(1+\tau)^{2}}\Big{|}_{\bm{\theta}_{t}}$
$\quad=\sum_{1\leq i<j\leq K}\frac{-2\tau}{(1+\tau)^{2}}\Big{(}r(x,y_{i})-r(x,y_{j})\Big{)}\Big{(}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x)|_{\bm{\theta}_{t}}-\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{j}|x)|_{\bm{\theta}_{t}}\Big{)}$
$\quad=\sum_{1\leq i<j\leq K}\frac{-2\tau}{(1+\tau)^{2}}\bigg{(}\big{(}r(x,y_{i})-r(x,y_{j})\big{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x)|_{\bm{\theta}_{t}}+\big{(}r(x,y_{j})-r(x,y_{i})\big{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{j}|x)|_{\bm{\theta}_{t}}\bigg{)}$
$\quad=\frac{-2\tau}{(1+\tau)^{2}}\sum_{1\leq i\leq K}\sum_{1\leq j\leq K}\big{(}r(x,y_{i})-r(x,y_{j})\big{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x)|_{\bm{\theta}_{t}}$
$\quad=\frac{-2\tau}{(1+\tau)^{2}}\sum_{1\leq i\leq K}K\cdot\big{(}r(x,y_{i})-\overline{r}(x)\big{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x)|_{\bm{\theta}_{t}},$
where $\overline{r}(x)\coloneqq\frac{1}{K}\sum_{j\in[K]}r(x,y_{j})$ in the last line.
Algorithm.
To this end, we update from $\bm{\theta}_{t}$ to $\bm{\theta}_{t+1}$ by taking one gradient step of the following surrogate loss, where we simplify the constant factor from $2\tau/(1+\tau)^{2}$ to $1/(1+\tau)$ and also drop the $K$ factor:
$\min_{\bm{\theta}}\quad\widehat{J}(\bm{\theta};x)\coloneqq-\frac{1}{1+\tau}\sum_{1\leq i\leq K}\big{(}r(x,y_{i})-\overline{r}(x)\big{)}\log\pi_{\bm{\theta}}(y_{i}|x).$
This is exactly the standard policy gradient using the group mean reward as the baseline, but derived differently with an off-policy interpretation. The hyperparameter $\tau\geq 0$ controls the size of each policy update. A larger $\tau$ is effectively the same as a smaller learning rate.
As a heuristic, we simply add a regularization term (denoted by $g$) to the above objective when additional regularization with respect to a fixed policy, e.g., a SFT model $\pi_{\textsf{sft}}$, is desired:
$\min_{\bm{\theta}}\quad\widehat{J}(\bm{\theta};x)\coloneqq-\frac{1}{1+\tau}\sum_{1\leq i\leq K}\big{(}r(x,y_{i})-\overline{r}(x)\big{)}\log\pi_{\bm{\theta}}(y_{i}|x)+\beta\cdot g\big{(}\pi_{\bm{\theta}},\pi_{\textsf{sft}};x,y_{1},\dots,y_{K}\big{)}.$
Appendix B Additional Details for Data Pipelines
B.1 Illustration of Typical Dataset Classes and Schemas
The dataset shaping process described in Section 3.3 is supported by PostgreSQL persistence with comprehensive schema designs; one example can be found in Listing 6. Key model classes include RftDatasetModel (core dataset storage with lineage tracking), ExperienceModel (which stores serialized experiences with priority scoring), DPODataModel and SFTDataModel (specialized storage for DPO and SFT, respectively).
1-- RftDataset table structure
2CREATE TABLE rft_dataset (
3 id SERIAL PRIMARY KEY, -- Unique record identifier
4 consumed_cnt INTEGER DEFAULT 0, -- Usage counter for sampling
5 prompt TEXT, -- Input instruction
6 response TEXT, -- Model-generated response
7 reward FLOAT, -- Training signal value
8 quality_score FLOAT DEFAULT 0.0, -- Quality assessment metric
9 difficulty_score FLOAT DEFAULT 0.0, -- Complexity estimation
10 diversity_score FLOAT DEFAULT 0.0, -- Response uniqueness measure
11 priority FLOAT DEFAULT 0.0 -- Sampling preference weight
12);
Listing 6: Schema definition for RftDataset PostgreSQL table.
B.2 Illustration of Data Processing Service
The architecture presented in Section 3.3 integrates Data-Juicer capabilities through service-oriented design. Typically, the DJ-service lifecycle covers three stages: initialization via scripts/start_servers.py; configuration through dj_config_path or natural language (dj_process_desc); and execution handling dataset loading, cleaning, priority scoring, and database export. Moreover, we provide DJExecutor to bridge Data-Juicer capabilities through config translation and execution sandboxing, and dynamic termination via _check_clean_targets() with melt protection.
B.3 An Example of Active Data Optimization Algorithms
The DataActiveIterator serves as the core engine for data-model co-evolution with multi-dimensional utility scoring. It provides extensible interfaces for flexible training prior injection, e.g., as demonstrated in Algorithm 1.
Algorithm 1 Multi-Dimensional Sample Selection
0:  Dataset $D$, weights $w_{q},w_{d},w_{k},w_{f}$
0:  Selected subset $D^{\prime\prime}$
1:  $(q,d,k)\leftarrow\text{DataCleaner}.\text{Process}(D)$ Compute quality ($q$), diversity ($d$), difficulty ($k$)
2:  $U_{i}\leftarrow w_{q}q_{i}+w_{d}d_{i}+w_{k}k_{i}-w_{f}f_{i}$ Combine $q$, $d$, $k$, frequency ($f$)
3:  $D^{\prime}\leftarrow\{s_{i}\in D\mid U_{i}>\theta_{\text{min}}\}$ Threshold filtering
4:  $D^{\prime\prime}\leftarrow\text{Top-K}(D^{\prime},K)$ Priority selection
5:  $\text{ExperienceCache}.\text{Update}(D^{\prime\prime})$ Update version-controlled cache
6:  $\text{LineageTracker}.\text{Record}(D^{\prime\prime})$ Track data provenance
7:  return $D^{\prime\prime}$

### v2

@@
 Reinforcement Fine-Tuning of Large Language Models
-[Uncaptioned image]: A General-Purpose and Unified Framework for
-Reinforcement Fine-Tuning of Large Language Models
-Xuchen Pan, Yanxi Chen, Yushuo Chen, Yuchang Sun, Daoyuan Chen,
-Wenhao Zhang, Yuexiang Xie, Yilun Huang, Yilei Zhang, Dawei Gao,
-Yaliang Li, Bolin Ding, Jingren Zhou
-Alibaba Group
-11footnotetext: Equal contribution.22footnotetext: Corresponding author. {chenyanxi.cyx, panxuchen.pxc, yaliang.li, bolin.ding}@alibaba-inc.com
-Trinity-RFT is a general-purpose, flexible and scalable framework designed for reinforcement fine-tuning (RFT) of large language models.
-It is built with a decoupled design, consisting of (1) an RFT-core that unifies and generalizes synchronous/asynchronous, on-policy/off-policy, and online/offline modes of RFT, (2) seamless integration for agent-environment interaction with high efficiency and robustness, and (3) systematic data pipelines optimized for RFT.
-Trinity-RFT can be easily adapted for diverse application scenarios, and serves as a unified platform for exploring advanced reinforcement learning paradigms.
-This technical report outlines the vision, features, design and implementations of Trinity-RFT, accompanied by extensive examples demonstrating the utility and user-friendliness of the proposed framework.
+Abstract
+Trinity-RFT is a general-purpose, unified and easy-to-use framework designed for reinforcement fine-tuning (RFT) of large language models.
+It is built with a modular and decoupled design, consisting of (1) an RFT-core that unifies and generalizes synchronous/asynchronous, on-policy/off-policy, and online/offline modes of RFT; (2) seamless integration for agent-environment interaction with high efficiency and robustness; and (3) systematic data pipelines optimized for RFT.
+Trinity-RFT can be easily adapted for diverse application scenarios, and serves as a unified platform for development and research of advanced reinforcement learning paradigms at both macroscopic and microscopic levels.
+This technical report outlines the vision, features, design and implementations of Trinity-RFT, accompanied by extensive examples, applications and experiments that demonstrate its functionalities and user-friendliness.
+11footnotetext: Equal contribution.22footnotetext: Corresponding author. {chenyanxi.cyx,panxuchen.pxc,daoyuanchen.cdy,yaliang.li,bolin.ding}@alibaba-inc.com
 Note: Trinity-RFT is currently under active development.
-This technical report corresponds to commit id f17db3d (May 23, 2025) of the GitHub repository, and will be continuously updated as the codebase evolves.
+This technical report (arXiv v2 version) corresponds to commit id 63d4920 (July 14, 2025) of the GitHub repository, and will be continuously updated as the codebase evolves.
 Comments, suggestions and contributions are welcome!
-Figure 1: The design of Trinity-RFT.
-1 Vision of Trinity-RFT
+1 Introduction
 Reinforcement learning (RL) has achieved remarkable success in the development of large language models (LLMs).
-Examples include aligning LLMs with human preferences via reinforcement learning from human feedback (RLHF) [18], and training long-CoT reasoning models via RL with rule-based rewards [3, 28].
-However, such approaches are limited in their abilities to handle dynamic, real-world, and continuous learning.
-Trinity-RFT envisions a future where AI agents learn by interacting directly with environments, collecting delayed or complex reward signals, and continuously refining their behavior through RL based on the collected experiences [26].
-For example, imagine an AI scientist who designs an experiment, executes it, waits for feedback (while working on other tasks concurrently), and iteratively updates itself based on true environmental rewards when the experiment is finally finished.
-Trinity-RFT offers a path into this future by providing various useful features.
-2 Key Features
-Trinity-RFT is a general-purpose, flexible, scalable and user-friendly RL framework that can be adapted for diverse experimental or real-world scenarios.
-The following features of Trinity-RFT make it a promising solution for realizing the aforementioned vision:
-  • Unified RFT modes & algorithm support.
-Trinity-RFT unifies and generalizes existing RFT methodologies into a flexible and configurable framework, supporting synchronous/asynchronous, on-policy/off-policy, and online/offline training, as well as hybrid modes that seamlessly combine them into a single learning process (e.g., incorporating expert trajectories to accelerate an online RL process [15, 36]).
-This is made possible partly by our decoupled design that allows rollout and training to be executed separately and scaled up independently on different machines, which will soon be introduced in the next section.
+Examples include aligning LLMs with human preferences via reinforcement learning from human feedback (RLHF) [23], and training long-CoT reasoning models via RL with rule-based or verifiable rewards (RLVR) [5, 34].
+However, such approaches are limited in their abilities to handle dynamic, agentic and continuous learning in the real world.
+We envision a future where AI agents learn by interacting directly with environments, collecting lagged or complex reward signals, and continuously refining their behavior through RL based on the collected experiences [31].
+For example, imagine an AI scientist who designs an experiment, executes it, waits for feedback (while working on other tasks concurrently), and iteratively updates itself based on true environmental rewards and feedback when the experiment is finally finished.
+This vision motivates us to develop Trinity-RFT, a reinforcement fine-tuning (RFT) framework that aims to offer a path into this future.
+The modular, decoupled and trinity design of Trinity-RFT illustrated in Figure 1, along with its various features, makes it a promising solution for realizing such a vision.
+Figure 1: The high-level design of Trinity-RFT.
+1.1 Key Features of Trinity-RFT
+Trinity-RFT is a general-purpose, unified, scalable and user-friendly RL framework that can be easily adapted for diverse experimental or real-world scenarios.
+It integrates both macroscopic and microscopic RL methodologies in one place; roughly speaking, the former deals with natural language and plain text, while the latter handles torch.Tensor (such as token probabilities, gradients, and model weights of LLMs) [^1] Many prior RL works for games/control/LLMs focus mainly on the microscopic aspect, e.g., designing policy loss functions or optimization techniques for updating the policy model.
+On the other hand, pre-trained LLMs, as generative models with rich prior knowledge of natural language and the world, open up numerous opportunities at the macroscopic level, e.g., experience synthesis by reflection or reasoning with environmental feedback [4], leveraging existing text processing methods like deduplication and quality filtering [2], among others..
+The key features of Trinity-RFT are presented below, which will be further elaborated in Section 2 and exemplified in Section 3.
+An RFT-core that unifies and generalizes diverse RL modes.
+Trinity-RFT implements diverse RL methodologies in a unified manner, supporting synchronous/asynchronous, on-policy/off-policy, and online/offline training.
+These RL modes can be flexibly generalized, e.g., a hybrid mode that incorporates expert trajectories to accelerate an online RL process [20, 45].
+This unification is made possible partly by our decoupled design (which will soon be introduced in Section 2.1) that allows rollout and training to be executed separately and scaled up independently on different devices, while having access to the same stand-alone experience buffer.
+The efficacy of various RL modes has been validated empirically by our experiments in Section 3.3, which particularly highlight the efficiency gains by off-policy or asynchronous methods.
 Agent-environment interaction as a first-class citizen.
-Trinity-RFT allows delayed rewards in multi-step/time-lagged feedback loops, and handles long-tailed latencies and environment/agent failures gracefully, which ensures efficiency and robustness in real-world scenarios with complex agent-environment interaction.
+Trinity-RFT allows delayed rewards and environmental feedback in multi-step/time-lagged feedback loop, handles long-tailed latencies and the straggler effect via asynchronous and streaming LLM inference, and deals with environment/agent failures gracefully via dedicated timeout/retry/skip mechanisms.
+These together ensure efficiency and robustness of continuous agent-environment interaction in complex real-world scenarios.
 Systematic data pipelines optimized for RFT.
-These include converting raw datasets to task sets for RFT, cleaning/filtering/prioritizing experiences stored in the buffer, incorporating priors and additional reward signals from various sources into the collected experiences, synthesizing data for augmenting tasks and experiences, offering user interfaces for human-in-the-loop, among others.
+Figure 2 illustrates the high-level design of data pipelines in Trinity-RFT, which regard rollout tasks and experiences as dynamic assets to be actively managed throughout the RFT lifecycle.
+Trinity-RFT empowers users to: (1) curate tasks for curriculum learning, e.g., by prioritizing easier tasks at the beginning of training to stabilize and accelerate the learning process; (2) actively manipulate experience by cleaning, filtering, or synthesizing new experiences, such as repairing failed trajectories or amplifying successful ones; (3) perform online reward shaping by augmenting sparse environmental rewards with dense, computed signals, such as quality or diversity scores; (4) customize interfaces for human-in-the-loop curation and utilize an agentic paradigm for RFT data processing that translates high-level natural language commands (e.g., “improve response diversity and safety for coding scenarios”) into complex data pipelines, powered by established community tools like Data-Juicer [2].
+For instance, Section 3.4 presents experiments that demonstrate the efficacy of task prioritization and reward shaping empowered by Trinity-RFT.
+Figure 2: The high-level design of data pipelines in Trinity-RFT.
+User-friendliness as a top priority.
+For development and research, the modular and decoupled design of Trinity-RFT allows the user to develop new RFT methodologies by adding one or a few small, plug-and-play classes (modified from built-in templates) that implement the essential functionalities of interest, with minimal code duplication or intrusive changes to the codebase.
+An example can be found in Section 3.2, which shows that three compact python classes (with around 200 lines of code in total) suffice for implementing a hybrid RL process that leverages samples from multiple data sources and updates the policy model with a customized loss function.
+For applications, the user can adapt Trinity-RFT to a new scenario by simply implementing a single Workflow class that specifies the logic of agent-environment interaction, as will be exemplified in Section 3.1.
+To further enhance usability, Trinity-RFT incorporates various graphical user interfaces to support low-code usage and development, enhance transparency of the RFT process, and facilitate easy monitoring and tracking.
+1.2 Related Works
+There exist numerous open-source RLHF frameworks, such as veRL [29], OpenRLHF [12], TRL [39], ChatLearn [1], Asynchronous RLHF [22], among others.
+Some of them have been further adapted for training long-CoT reasoning models or for agentic RL more recently.
+Concurrent to Trinity-RFT, some recent works on LLM reinforcement learning also advocate a decoupled and/or asynchronous design; examples include StreamRL [49], MiMo [43], AReaL [8], ROLL [36], LlamaRL  [42], Magistral [17], AsyncFlow [11], among others.
+Complementary to this huge number of related works, Trinity-RFT provides the community with a new solution that is powerful, easy-to-use, and unique in certain aspects.
+In a nutshell, Trinity-RFT aims to be general-purpose and applicable to diverse application scenarios, while unifying various RFT modes, RFT methodologies at macroscopic and microscopic levels, and RFT-core/agent-environment interaction/data pipelines.
+Such a system-engineering perspective makes Trinity-RFT particularly useful for handling the whole RFT pipeline in one place.
+We also hope that some specific features of Trinity-RFT, such as data persistence in the experience buffer, and distributed deployment of multiple independent explorers, will open up new opportunities for LLM reinforcement fine-tuning.
 2 Design and Implementations
@@
 RFT-core is the component of Trinity-RFT, highlighted at the center of Figure 1, where the core RFT process happens.
-Its design also exhibits a trinity, consisting of the explorer, trainer, and buffer.
+Its design also exhibits a trinity, consisting of the explorer, buffer, and trainer.
   • The explorer, powered by a rollout model, takes a task as input and solves it by executing a workflow that specifies the logic of agent-environment interaction, thereby collecting experiences (including rollout trajectories, rewards, and other useful information) to be stored in the buffer.
   • The buffer stores experiences that can be generated by the explorer or by other sources, such as human experts.
+It can be realized in various forms, such as a non-persistent ray.Queue or a persistent SQLite database.
 It also assists with fetching training samples for the trainer, and can be integrated with advanced sampling strategies and post-processing operations.
   • The trainer, backed by a policy model, samples batches of experiences from the buffer and updates the policy model via RL algorithms.
-Our implementations allow the explorer and trainer to be deployed on separate machines.
-This decoupled design of RFT-core enables Trinity-RFT to support diverse RFT modes with great flexibility, for example:
-  • In a synchronous mode, the explorer and trainer get launched simultaneously, work in close coordination, and synchronize their model weights once every sync_interval training steps.
+Our implementations allow the explorer and trainer to be deployed on separate machines and act independently.
+They are only connected via (1) access to the same experience buffer with a customizable sampling strategy, and (2) model weight synchronization by a customizable schedule.
+See Figure 3 for an illustration.
+This decoupled design of RFT-core offers support for diverse RFT modes with great flexibility.
+Figure 3: The architecture of RFT-core in Trinity-RFT.
+2.1.1 Unified Support for Diverse RFT Modes
+We present the RFT modes supported by Trinity-RFT, some of which are demonstrated in Figure 4.
+Figure 4: A visualization of diverse RFT modes supported by Trinity-RFT, including: (a) synchronous mode, with sync_interval=2; (b) one-step off-policy mode, with sync_interval=1 and sync_offset=1; (c) fully asynchronous mode, with sync_interval=2; (d) multi-explorer asynchronous mode, with sync_interval=2.
+The buffer supports, in principle, arbitrary management and sampling strategies for experiences.
+Synchronous mode.
+In the synchronous mode shown in Figure 4 (a), the explorer and trainer get launched simultaneously, work in close coordination, and synchronize their model weights once every sync_interval training steps.
 Within each synchronization period, the explorer continuously generates sync_interval batches of rollout experiences and stores them in the buffer, which are then retrieved and utilized by the trainer for updating the policy model.
-If sync_interval $=1$, this is an on-policy RL process, whereas if sync_interval $>1$, it becomes off-policy (akin to the mode adopted in [28]) and can be accelerated by pipeline parallelism between the explorer and trainer.
-This mode can be activated by setting the configuration parameter mode to both.
-  • In a fully asynchronous mode, the explorer and trainer are mostly decoupled and can scale up independently.
+If sync_interval=1, this is a strictly on-policy RL process, whereas if sync_interval>1, it becomes off-policy (akin to the mode adopted in [34]) and can be accelerated by pipeline parallelism between the explorer and trainer.
+This mode can be activated by setting the configuration parameter mode=both.
+One-step off-policy mode.
+This mode, demonstrated in Figure 4 (b), closely resembles the synchronous mode, except for an offset of one batch between the explorer and trainer.
+This allows the trainer to sample experiences from the buffer immediately after model weight synchronization, thereby streamlining the execution of explorer and trainer with smaller pipeline bubbles, at the cost of slight off-policyness.
+The visualization in Figure 4 (b) corresponds to configuration parameters sync_interval=1 and sync_offset=1, both of which can take more general values in Trinity-RFT.
+Asynchronous mode.
+In the fully asynchronous mode shown in Figure 4 (c), the explorer and trainer act almost independently.
 The explorer continuously generates rollout experiences and stores them in the buffer, while the trainer continuously samples experiences from the buffer and uses them for training the policy model.
 External experiences, e.g., those generated by expert models or humans, can be continuously incorporated into the buffer as well.
-The explorer synchronizes its model weights with the trainer once in a while, keeping the distribution of rollout experiences up to date.
-This mode can be activated by setting mode to explore/train and launching the explorer and trainer separately on different machines.
+The explorer and trainer independently load or save model weights from the checkpoint directory every sync_interval steps, keeping the distribution of rollout experiences up to date.
+This mode can be activated by setting mode=explore/train and launching the explorer and trainer separately on different GPUs.
+Multi-explorer asynchronous mode.
+One benefit brought by the decoupled design is that explorers and trainers can scale up independently on separate devices.
+As a proof-of-concept, Trinity-RFT offers support for a multi-explorer asynchronous mode, as demonstrated in Figure 4 (d), where multiple explores send the generated rollout experiences to the same buffer.
+Scaling up the number of independent and distributed explorers can be particularly useful for resolving data scarcity and speeding up the generation of experiences in real-world scenarios where rollout trajectories have to be sampled via interaction with the physical world, or in an environment with sparse and lagged feedback.
+Another by-product of this multi-explorer mode is 24-hour non-interrupted service for real-world online serving situations: since the explorers can pause and update model weights at different moments, it can be guaranteed that there is always one explorer ready to serve an incoming request immediately whenever it arrives.
+This is in contrast to a single-explorer mode, where online service has to be paused when the explorer is updating its model weights.
+Benchmark mode.
 Trinity-RFT supports a benchmark mode that allows the user to evaluate one or multiple checkpoints on arbitrary benchmarks, after the RFT training process has finished.
-To activate this mode, the user simply needs to set mode to bench and specify the paths for the evaluation datasets in the configurations.
+To activate this mode, the user simply needs to set mode=bench and specify the paths for the evaluation datasets in the configurations.
 This mode can be particularly useful for experimental purposes; for example, the user might want to try out different RFT techniques or configurations quickly (with limited evaluation on hold-out data) during training, identify which RFT trials have achieved stable convergence and high rewards, and then conduct more thorough evaluations only for the checkpoints of these successful trials.
+Train-only mode.
 In certain scenarios, the user would like to train the policy model without further exploration, using experiences that have already been collected and stored in the buffer.
-This train-only mode can be activated by setting the mode parameter to train and launching the trainer alone.
+This train-only mode can be activated by setting the configuration parameter mode $=$ train and launching the trainer alone.
 Offline methods like Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) [24] can be regarded as special cases of such scenarios, both of which are natively supported by Trinity-RFT.
-For another example, consider an online RFT process that expands over a long period of time, where the explorer alone is launched during the daytime for serving human users and collecting experiences, while the trainer alone is launched at night for updating the policy model (which will be thoroughly validated and evaluated before it can be actually deployed as the rollout model for the next day).
-Another benefit brought by the decoupled design is that explorers and trainers can operate across separate devices (e.g., multiple explorers on edge devices, one trainer on a high-end GPU cluster) and scale up independently.
-Indeed, we anticipate that for successful RL in real-world scenarios, distributed collection of massive rollout experiences could be one promising solution to resolving data scarcity.
-Remark 1.
-We take a system-algorithm co-design perspective in the development of Trinity-RFT, and aim to unify and generalize diverse RFT methodologies in this framework, including synchronous/asynchronous, on-policy/off-policy, and online/offline modes.
+For another example, consider an online RFT process that expands over a long period, where the explorer alone is launched during the daytime for serving human users and collecting experiences, while the trainer alone is launched at night for updating the policy model (which will be thoroughly validated and evaluated before it can be actually deployed as the rollout model for the next day).
+Discussions.
+We conclude this subsection with two remarks.
+(1) Given the unified implementation of various RFT modes, it is easy to design and implement a hybrid mode with Trinity-RFT that combines multiple modes into a single learning process.
+One example is learning with both online rollout data and offline-collected expert data, via jointly optimizing two loss terms corresponding to these two data sources.
+Section 3.2 illustrates how to implement this conveniently in Trinity-RFT.
+(2) We take a system-algorithm co-design perspective in the development of Trinity-RFT, aiming to unify and generalize diverse RFT methodologies in this framework.
 RFT-core provides the necessary infrastructure for achieving this goal.
 While the main text of this technical report focuses on the system perspective, interested readers may refer to Appendix A for further discussions on the algorithm perspective of off-policy/asynchronous RL.
+2.1.2 Implementations of RFT-Core
+We present some implementation details of RFT-core in the following.
+Inference and training engines.
+The current version of Trinity-RFT leverages vLLM [14] as the inference engine for the explorer, which offers features including paged attention, continuous batching [48], asynchronous and concurrent inference for multiple rollout trajectories, among others.
+Trinity-RFT also leverages verl [29] as the training engine for the trainer, which gracefully handles model placement (for the policy, critic and reference models) and incorporates various performance optimizations for training (such as dynamic batching, management of padding and unpadding, etc.).
+Trinity-RFT stands on the shoulders of these excellent open-source projects, and will continue to benefit from their future development.
+Experience buffer.
+Trinity-RFT supports multiple types of experience buffers, ranging from a non-persistent ray.Queue to persistent SQLite or Redis databases.
+While using a basic first-in-first-out queue is the most straightforward approach, data persistence with a database opens up many new opportunities (e.g., advanced sampling strategies), as discussed throughout this report.
+Trinity-RFT has provided dedicated read/write control to prevent any conflict in accessing the buffer.
+Model weight synchronization.
+Trinity-RFT supports model weight synchronization between the explorer and trainer by NCCL [21], or by checkpoint saving and loading.
+The former is faster (when available), while the latter is generally more flexible and widely applicable, especially for asynchronous RFT modes.
 2.2 Agent-Environment Interaction
+To adapt Trinity-RFT to a new downstream scenario, the user mainly needs to define and register a customized workflow (by inheriting the base class Workflow or MultiTurnWorkflow) where the logic of agent-environment interaction for this particular scenario is implemented.
+Advanced methods for experience synthesis with environmental feedback [4] can be implemented in the same way as well.
+See Section 3.1 for detailed examples.
+The workflow will then be executed by workflow runners within the explorer for generating experiences, as shown in Figure 3.
 Numerous challenges arise when one tries to build an RFT framework that can efficiently and robustly handle real-world interaction between the LLM-powered agent and the environment.
@@
 Trinity-RFT regards agent-environment interaction as a first-class citizen and incorporates various solutions to tackle these challenges, for example:
-  • Built upon vLLM [10], our implementations of the LLM inference engine in Trinity-RFT support asynchronous and concurrent generation of rollout trajectories for multiple tasks.
+  • The workflow runners in Trinity-RFT support asynchronous and streaming generation of rollout trajectories for multiple tasks.
 This helps mitigate the straggler effect caused by the long-tailed latencies in rollout generation and agent-environment interaction, thereby accelerating the RFT process.
@@
 When the reward signals from the environment finally arrive, they are written to the buffer, and the corresponding experience is now marked as “ready for training”.
-To adapt Trinity-RFT to a new downstream scenario, the user mainly needs to define and register a customized workflow (by inheriting the base class Workflow or MultiTurnWorkflow) where the logic of agent-environment interaction is implemented.
-Detailed examples will be demonstrated in Sections 4.1 and 4.2.
+  • For multi-turn conversations and ReAct-style workflows [47], Trinity-RFT supports concatenating multiple rounds of agent-environment interaction compactly into a single sequence, with proper masking that indicates which tokens need to be incorporated into the training objective of RL algorithms.
+This avoids unnecessary recomputation and thus improves training efficiency, compared to a vanilla approach that represents a $K$-turn rollout trajectory with $K$ separate samples.
+  • As another performance optimization, the implementation of Trinity-RFT allows resetting the environment in a workflow, rather than re-initializing it every time.
+This is especially useful for scenarios where setting up the environment is costly.
 2.3 Data Pipelines
 The data pipelines in Trinity-RFT aim to address fundamental challenges in RFT scenarios, such as managing heterogeneous data dynamics across interaction workflows, enabling delayed reward integration, and facilitating continuous data curation.
-Our solutions center on three core aspects: end-to-end data transformation, active data-reward co-shaping, and human-in-the-loop curation, each corresponding to key requirements identified in our development of RFT-core (Section 3.1).
+Our solutions center on four core aspects: end-to-end data transformation, task curation, active experience shaping, and human-in-the-loop curation, each corresponding to key requirements identified in our development of RFT-core (Section 2.1).
 2.3.1 End-to-end Data Transformation
@@
 It decouples data pipeline logic from procedure control to enable flexible RL-oriented data transformations with two key modules:
-  • The Formatter Module unifies disparate data sources into RFT-compatible formats, providing convenient conversion between raw inputs (e.g., meta-prompts, domain-specific corpora, and QA pairs with tagged rewards) and structured RFT representations (e.g., via RftDatasetModel).
-For efficient RFT workloads, we utilize PostgreSQL-based persistent storage supporting both ExperienceModel for prioritized trajectories and DPODataModel for preference pairs, which is particularly beneficial for asynchronous RFT mode and distributed environments.
-Full schema details appear in Appendix B.1.
+  • The Formatter Module unifies disparate data sources into RFT-compatible formats, providing convenient conversion between raw inputs (e.g., meta-prompts, domain-specific corpora, and QA pairs with tagged rewards) and structured RFT representations.
+For efficient RFT workloads, we utilize buffer-based persistent storage (Section 2.1) to support different data models, such as ExperienceModel for prioritized rollout trajectories and DPODataModel for preference pairs.
+The conversion logic and data models are highly customizable to meet diverse requirements for managing experience data.
+This flexibility enables robust metadata recording and field normalization, which is essential for advanced scenarios such as asynchronous RFT in trainer-explorer environments, agent self-evolution from a cold start using meta-prompts, and knowledge injection from structurally complex domain-specific corpora.
   • The Controller Module manages the complete data pipeline lifecycle through distributed server initialization, declarative configuration, and automated dataset persistence.
@@
 This modular design enables Trinity-RFT to handle data transformations flexibly while maintaining consistency across different RFT modes.
-More service utilities are detailed in Appendix B.2.
 The Formatter-Controller duality mirrors the explorer-trainer decoupling in RFT-core, enabling parallel data ingestion and model updating.
 This design also allows Trinity-RFT to handle delayed rewards through version-controlled experience updates while maintaining low-latency sampling for the trainer.
-Figure 2: An illustration of data flow in Trinity-RFT.
-3.3.2 Active Data-Reward Co-Shaping
-To enable efficient and effective learning from complex and delayed reward signals, Trinity-RFT introduces fruitful mechanisms that actively shape both data and reward throughout the RFT process.
-  • Trinity-RFT provides the Processor Module to support convenient and extensive data processing capabilities tailored to RFT, through dedicated bridge classes that link Trinity-RFT to over 100 operators from the Data-Juicer system [2].
-This module offers efficient integration with the experience buffer, and provides composable building blocks for experience cleaning (e.g., length filters, duplication removal), safety alignment (e.g., toxicity detection, ethics checks), and preference data synthesis (e.g., critique-conditioned augmentation).
-Users can specify processing chains using either natural language or explicit configurations.
-These are accessible to both RFT users familiar with Data-Juicer and those who are not, thanks to transparency of the underlying system details.
+Figure 5: The interaction of data processor and data buffers in Trinity-RFT, divided into two key stages.
+Left: Task Curation & Prioritization prepares the initial tasks for the explorer.
+Right: Experience Shaping processes the collected trajectories from the explorer before they are used by the trainer.
+The data processor is a central component that operates on different buffers at different stages.
+2.3.2 Task Curation and Prioritization
+Before the RFT loop begins, it is crucial to prepare a high-quality set of initial tasks.
+This stage, depicted on the left side of Figure 5, transforms raw data into an optimized task set for the explorer.
+The process begins with raw data sources (e.g., prompts, domain corpora), which are ingested into a buffer.
+The Data Processor, powered by over 100 operators from Data-Juicer [2], reads from this buffer to perform various curation tasks.
+It provides composable building blocks for experience cleaning (e.g., length filters, duplication removal), safety alignment (e.g., toxicity detection, ethics checks), and preference data synthesis (e.g., critique-conditioned augmentation).
 By treating Data-Juicer as a modular data processing operator pool rather than a central dependency, Trinity-RFT provides RL-specific abstractions and coherence, while benefiting from well-established data tools.
-  • RL-oriented data shaping.
-RL is particularly useful in scenarios where optimization goals cannot be clearly defined, leading users to express vague or high-level requirements.
-In response, Trinity-RFT automates the translation of these objectives into optimized data processing pipelines.
-It reduces problem complexity through a structured hierarchy: Quality $\rightarrow$ Difficulty $\rightarrow$ Diversity $\rightarrow$ Quantity, with auto-selection of appropriate operators from Data-Juicer.
-Moreover, Trinity-RFT provides many composable and extensible interfaces for data shaping, aiming to seamlessly inject user-defined inductive biases.
-For instance, DataCleaner (a sub-module of Processor) supports multi-way cleaning with statistical filtering (e.g., entropy-based diversity [11]) and rule-based validation (e.g., boxed answer verification in math dataset [7]), while DataSynthesizer enables both autonomous generation and hybrid augmentation with controlled feedback metrics [35].
-  • Active data optimization.
-As agents interact with environments, experiences collected and stored in the buffer can become large-scale and highly diverse.
-Using the right experiences for RL training is crucial for a cost-effective and stable learning process.
-To this end, Trinity-RFT allows users to flexibly apply multi-dimensional utility scoring to prioritize samples (e.g., 0.4 * difficulty + 0.3 * diversity + 0.1 * quality - 0.2 * frequency).
-Moreover, we introduce the DataActiveIterator module to enable (1) version-controlled experience reuse through ExperienceModel; (2) cross-task data lineage tracking via rich relationships derived from sample_id and rollout-model_id; and (3) incorporating human annotations into the rewards.
-These features will be demonstrated through concrete examples in Section 4 and Appendix B.3.
+The processed data is then organized into a structured task buffer.
+This stage effectively implements a form of curriculum learning by allowing users to prioritize tasks (e.g., from easy to hard), guiding the explorer towards a more efficient and stable learning trajectory from the outset.
+This entire workflow is managed by a service-oriented architecture that decouples data logic from procedural control, ensuring flexibility and scalability, especially in asynchronous and distributed settings.
+2.3.3 Active Experience Shaping
+Once the explorer begins interacting with the environment, it generates a continuous stream of experience data.
+To maximize learning efficiency, this raw experience must be actively shaped before it reaches the trainer.
+This stage is shown on the right side of Figure 5.
+Generated experiences are first collected in a buffer.
+The Data Processor is applied again with a series of transformations to clean, augment, or synthesize these experiences.
+This is where the core of RFT data intelligence lies.
+Key capabilities include:
+  • Agent-Driven Data Processing: Trinity-RFT introduces a powerful agentic paradigm for data manipulation.
+Users can define complex processing pipelines through high-level objectives, specified as either natural language commands (e.g., “improve safety” or “increase response diversity”) or explicit Data-Juicer configurations.
+The framework automatically translates these commands into executable workflows backed by its modular components like DataCleaner and DataSynthesizer.
+This design provides a user-friendly abstraction layer over the underlying Data-Juicer operators, making advanced processing functionalities accessible to both RFT users familiar with Data-Juicer and those who are not.
+It also facilitates the flexible injection of user-defined inductive biases into the learning process, unlocking new research directions for self-evolving agents, as we will discuss later in Section 2.3.5.
+  • Online Reward Shaping: The data processor can dynamically augment the reward signal.
+Instead of relying on a single, often sparse, task-completion reward, users can add dense rewards based on quality, diversity, or safety scores computed on the fly.
+This enriched feedback provides a much stronger learning signal for the trainer.
+  • Prioritized Experience Replay: Experiences are not treated equally.
+Trinity-RFT allows for flexible, multi-dimensional utility scoring to prioritize the most valuable samples for training.
+The DataActiveIterator supports version-controlled experience reuse and cross-task data lineage tracking, ensuring that the trainer always learns from the most informative data available.
+This mechanism is also critical for handling delayed rewards, as experience utilities can be updated asynchronously as new feedback arrives.
 2.3.4 Human-AI Collaboration
@@
 Moreover, for easy deployment, we provide local Label Studio instance management with automatic environment setup via Docker/pip; optimized SDK interactions with batch request coalescing; unified logging across annotation tools and ML services; and concurrent annotation campaigns through priority-based task routing, while maintaining full data lineage preserved via LineageTracker.
-The decoupled design of Trinity-RFT, or the presence of an experience buffer in particular, enables human feedback to participate in RL loops without breaking the asynchronous execution model.
+The decoupled design of Trinity-RFT, and the presence of a standalone experience buffer in particular, enable human feedback to participate in RL loops without breaking the asynchronous execution model.
 For instance, human-verified samples can be prioritized for training while fresh experiences are being collected, which is a critical capability for real-world deployment scenarios with mixed feedback sources.
 Further details for human-AI collaboration in Trinity-RFT will be illustrated in Section 3.5.
-3.4 Performance Optimizations and User-Friendliness
-Trinity-RFT has made numerous efforts in pursuit of efficiency, for example:
-  • For multi-turn conversations and ReAct-style workflows [37], Trinity-RFT supports concatenating multiple rounds of agent-environment interaction compactly into a single sequence, with proper masking that indicates which tokens need to be incorporated into the training objective of RL algorithms.
-  • For the synchronous RFT mode, Trinity-RFT accelerates the RFT process via pipeline parallelism between the explorer and trainer.
-  • Trinity-RFT utilizes Ray [14] for distributed runtime, and NCCL communication primitives [16] (whenever feasible) for model weight synchronization between the explorer and trainer.
-In addition, Trinity-RFT has been designed with user-friendliness as a top priority, for example:
-  • We include a monitor (built upon Wandb [32] and TensorBoard [29]) that allows the user to conveniently track the progress of an RFT process, both quantitatively (e.g., via learning curves for rewards and other metrics) and qualitatively (e.g., via concrete examples of rollout trajectories generated at different RL steps).
+2.3.5 Discussion: Unlocking New Research & Development Directions
+The modular design of our data pipelines and the powerful data processor open up promising research and development avenues to be further explored.
+One direction is about effective management of experience data.
+While prior RFT works often treat the experience as a static log, Trinity-RFT enables a more sophisticated, full-lifecycle approach to data, from selective acquisition to efficient representation:
+  • Intelligent Perception and Collection: In an open-ended environment, what experience is “worth” recording?
+Storing everything creates a low signal-to-noise ratio and burdens the trainer.
+Trinity-RFT’s architecture allows researchers to implement active collection strategies.
+For instance, one could design a data processor operator that evaluates incoming experiences from the explorer based on metrics like surprise, uncertainty, or information gain, and only commits the most salient trajectories to the replay buffer.
+This transforms data collection from passive logging into a targeted, intelligent process.
+  • Adaptive Representation: Raw experience is often high-dimensional and redundant (e.g., long dialogues, complex code generation traces).
+How can this be distilled into a format that an agent can efficiently learn from?
+The data processor in Trinity-RFT acts as a powerful transformation engine.
+Researchers can use it to explore various representation learning techniques, such as automatically summarizing trajectories, extracting causal chains from tool usage, or converting a multi-turn dialogue into a structured preference pair.
+This not only makes training more efficient but also opens the door to building meta-experience (more abstract and reusable knowledge) from raw interaction data.
+  • Agentic Workflows: Trinity-RFT’s agent-driven processing enables the research and development of self-improving agents, e.g., by configuring the policy agent to also serve as the “processing agent” for LLM-based Data-Juicer operators.
+Such an agent could perform its own critique and dynamically curate its own training data, creating a truly autonomous learning and data management loop.
+Another direction is about synthetic and counterfactual experience processing.
+The integration of synthesis operators enables research into creating “better-than-real” data.
+Instead of relying solely on the agent’s own trial-and-error, our framework facilitates exploring questions like:
+  • Dynamic and Composable Rewarding: With our framework, researchers can move beyond static, hand-crafted rewards.
+It is now possible to investigate dynamic reward shaping, where auxiliary signals like novelty, complexity, or alignment scores are automatically extracted from trajectories and composed into a dense reward function.
+How to define “good” experience and how can we learn the optimal combination of these reward components as the agent’s policy evolves?
+  • Experience Reorganization: Can successful sub-trajectories from different tasks be “spliced” together to solve a novel, composite task?
+For example, can an agent that has learned to “open a door” and “pick up a cup” synthesize a new trajectory to "enter the room and fetch the cup"?
+  • Failure Repair: Can the data processor identify where errors occur in a failed trajectory, and synthesize a corrected version for the trainer to learn from, effectively turning failures into valuable lessons?
+  • Success Amplification: Can a single successful experience be augmented into multiple diverse yet successful variants, thereby improving the generalization and robustness of the learned policy?
+By providing dedicated capabilities for such advanced data and reward manipulation, Trinity-RFT aims to facilitates flexible processing of “experience data” for the next generation of self-evolving LLMs.
+2.4 User-Friendliness
+Trinity-RFT has been designed with user-friendliness as a top priority.
+For development and research:
+The modular and decoupled design of Trinity-RFT allows users to develop a new algorithm for a specific aspect of RFT by adding one or a few new classes that implement the essential functionalities of interest, without concerning about other aspects of RFT or intrusive modifications of the original codebase.
+In addition, we include a monitor (built upon Wandb [40] and TensorBoard [37]) that makes it easy to track the progress of an RFT process, both quantitatively (e.g., via learning curves for rewards and other metrics) and qualitatively (e.g., via concrete examples of rollout trajectories generated at different RL steps).
 See Figure 6 for an example snapshot of the monitor.
-  • We also implement a configuration manager, as shown in Figure 4, that allows the user to create configuration files conveniently via a front-end interface.
-  • We provide Trinity-Studio, an all-in-one unified UI that allows the user to configure and run data inspection, data processing, RFT learning process, etc., all by clicking the mouse and filling forms, without writing any code.
-More functions will be integrated in the future.
-An example for developing with Trinity-Studio will be introduced in Section 4.5.
+For RFT applications:
+Trinity-RFT offers extensive graphical user interfaces to support low-code usage of the framework, and to maximize transparency of the RFT process.
+For example, we implement a configuration manager, as shown in Figure 7, that allows the user to create configuration files conveniently via a front-end interface.
+We also provide Trinity-Studio, an all-in-one unified UI (including the aforementioned monitor and configuration manager) that allows the user to configure and run data inspection, data processing, RFT learning process, etc., all by clicking the mouse and filling forms, without writing any code.
+An example for using Trinity-Studio will be introduced in Section 3.6.
+Such functionalities, of course, can be useful not only for applications but also for development and research.
 Figure 6: A snapshot of the monitor implemented in Trinity-RFT.
@@
 Figure 7: Snapshots of the configuration manager.
-4 Examples and Applications
-Through some concrete examples, we demonstrate how to use Trinity-RFT for different purposes, and also exemplifies some concepts introduced in previous sections.
-Additional step-by-step tutorials can be found on the documentation website [^1] https://modelscope.github.io/Trinity-RFT, or the /examples folder of the GitHub repository [^2] https://github.com/modelscope/Trinity-RFT.
-4.1 Single-Turn Workflow
+3 Examples, Applications, and Experiments
+This section demonstrates the utilities and user-friendliness of Trinity-RFT and exemplifies some concepts introduced in previous sections, through a diverse range of examples, applications and experiments.
+Additional step-by-step tutorials can be found on the documentation website [^2] https://modelscope.github.io/Trinity-RFT, or the examples folder of the GitHub repository [^3] https://github.com/modelscope/Trinity-RFT/tree/main/examples.
+3.1 Customizing Agent-Environment Interaction
+With a modular design, Trinity-RFT can be easily adapted to a new downstream scenario by implementing the logic of agent-environment interaction in a single workflow class, without modifications to other components of the codebase.
+This approach is also sufficient for macroscopic RL algorithm design that targets high-quality experience synthesis with environmental feedback [4].
+We provide some concrete examples in the rest of this subsection.
+3.1.1 Single-turn Scenarios
 In a simple yet common scenario, a user of Trinity-RFT would like to train an LLM for completing single-turn tasks, where the LLM generates one response to each input query.
-For this purpose, the user mainly needs to (1) define and register a single-turn workflow class (by inheriting the base class Workflow) tailored to the targeted tasks, and (2) specify the dataset of tasks (for training and/or evaluation) and the initial LLM, both compatible with HuggingFace [9] and ModelScope [13] formats.
+For this purpose, the user mainly needs to (1) define and register a single-turn workflow class (by inheriting the base class Workflow) tailored to the targeted tasks, and (2) specify the tasksets (for training and/or evaluation) and the initial LLM, both of which are compatible with HuggingFace [13] and ModelScope [18] formats.
 Listing 1 gives a minimal example for implementing a single-turn workflow.
 Suppose that each task is specified by a <question, answer> tuple.
-The run() method of ExampleWorkflow calls the LLM once to generate a response for the question, calculates its reward, and returns an Experience instance that consists of the response itself, the reward value, and the log-probabilities of next-token prediction by the rollout model (which is necessary for certain RL algorithms, such as PPO [22] and GRPO [23]).
-Some built-in workflows have been implemented in Trinity-RFT, e.g., the MathWorkflow class for math-related tasks.
-Trinity-RFT also allows more flexible and advanced usage by changing just a few configuration parameters, e.g., running supervised fine-tuning (SFT) as a warmup stage before the RFT process, or choosing between the synchronous and asynchronous modes of RFT (as described in Section 3.1).
+The run() method of ExampleWorkflow calls the LLM once to generate a response for the question, calculates its reward, and returns an Experience instance that consists of the response itself, the reward value, and the log-probabilities of response tokens predicted by the rollout model (which is necessary for certain RL algorithms, such as PPO [27] and GRPO [28]).
+Some built-in workflows and reward functions have been implemented in Trinity-RFT, e.g., the MathWorkflow class for math-related tasks.
+In some cases, the user wants to utilize auxiliary LLMs in the workflow, e.g., for computing rewards via LLM-as-a-judge, or for playing the roles of other agents in a multi-agent scenario.
+For these purposes, the user can specify auxiliary_models via APIs when initializing the workflow.
 1@WORKFLOWS.register_module("example_workflow")
@@
 3
-4 def __init__(self, model: ModelWrapper, task: Task, **kwargs):
-5 super().__init__(model, task, **kwargs)
+4 def __init__(
+5 self,
+6 model: ModelWrapper,
+7 task: Task,
+8 auxiliary_models: Optional[List[openai.OpenAI]] = None,
+9 ):
+10 super().__init__(model, task, auxiliary_models)
 11 self.question = task.raw_task.get("question")
 12 self.answer = task.raw_task.get("answer")
-8
-9 def calculate_reward(self, response: str, truth: str) -> float:
+13
+14 def calculate_reward_by_rule(self, response: str, truth: str) -> float:
 15 return 1.0 if response == truth else 0.0
-11
+16
+17 def calculate_reward_by_llm_judge(self, response: str, truth:str) -> float:
+18 judge_model = self.auxiliary_models[0]
+19 PROMPT_FOR_JUDGE = """Please evaluate..."""
+20 completion = judge_model.chat.completions.create(
+21 model="gpt-4", # Or another suitable judge model
+22 messages=[{"role": "user", "content": PROMPT_FOR_JUDGE}],
+23 )
+24 reward_str = completion.choices[0].message.content.strip()
+25 reward = float(reward_str)
+26 return reward
+27
 28 def run(self) -> List[Experience]:
@@
 35 ],
-20 n=self.task.rollout_args.repeat_times,
-21 temperature=self.task.rollout_args.temperature,
+36 **self.rollout_args,
 37 )
-23 reward: float = self.calculate_reward(response.response_text, self.answer)
+38 reward: float = self.calculate_reward_by_rule(response.response_text, self.answer)
+39 # reward: float = self.calculate_reward_by_llm_judge(response.response_text, self.answer)
 40 return [
@@
 Listing 1: A minimal example for implementing a customized workflow.
-4.2 Multi-Turn Workflow
+3.1.2 Multi-turn Scenarios
 In more advanced cases, the user would like to train an LLM-powered agent that solves multi-turn tasks by repeatedly interacting with the environment.
-With Trinity-RFT, achieving this is mostly as simple as in the single-turn case, except that the user needs to define and register a multi-turn workflow class (by inheriting the base class MultiTurnWorkflow), where the logic of agent-environment interaction is implemented.
-Listing 2 provides one such example for ALFWorld [25].
+In Trinity-RFT, achieving this is mostly as simple as in the single-turn case, except that the user needs to define and register a multi-turn workflow class by inheriting the base class MultiTurnWorkflow.
+Listing 2 provides one such example using the ALFWorld dataset [30].
 For training efficiency, the process_messages_to_experience() method concatenates multiple rounds of agent-environment interactions compactly into an Experience instance consisting of a single token sequence with proper masking, which can readily be consumed by standard RL algorithms like PPO and GRPO.
+For more detailed examples of multi-turn cases, please refer to the documentation [^4] https://modelscope.github.io/Trinity-RFT/tutorial/example_multi_turn.html.
 1@WORKFLOWS.register_module("alfworld_workflow")
@@
 4
-5 # ...
-6
 5 def generate_env_inference_samples(self, env, rollout_num) -> List[Experience]:
@@
 15 memory = memory + [{"role": "user", "content": format_obs}]
-18 response_text = self.get_model_response_text(memory)
+16 response_text = self.model.chat(memory, n=1)[0].response_text
 17 memory.append({"role": "assistant", "content": response_text})
@@
 29 return experience_list
-32
-33
+30
 31 def run(self) -> List[Experience]:
@@
 Listing 2: An implementation of a multi-turn workflow for ALFWorld [30].
-4.3 Dataset Curation
-We present an example to showcase the end-to-end data transformation (Section 3.3.1) and data-reward co-shaping (Section 3.3.2) capabilities of Trinity-RFT in a mathematical reasoning task.
-As shown in Listings 3 and 4, the user configures dataset paths and field mappings through a centralized DataConfig, then specifies natural-language processing instructions via dj_process_desc.
-The system will automatically translate these requirements into executable pipelines through three phases: (1) Processor applies LLM-powered difficulty scoring using Qwen-Max; (2) Controller optimizes quality-diversity tradeoffs through adaptive filtering thresholds and thereby prioritizes high-value training samples; (3) Formatter converts raw arithmetic problems into an RL-ready task set with specified reward functions and workflows.
-This procedure exemplifies our active data optimization capability (Section 3.3.2) by combining automated quality assessments with configurable curation policies.
-Users can extend this pattern to many other domains through customization and adaptation: for example, users can define their own difficulty metrics and calculation rules in the parameters of Data-Juicer’s OPs.
-Besides, users can flexibly utilize another Synthesizer module and extend new Controller policies to support seamless transitions between different domain-specific reward shaping and data optimization priors.
+3.1.3 Experience Synthesis in Workflows
+As mentioned in Section 1.1, Trinity-RFT has been designed to streamline RL algorithm design and development at both macroscopic and microscopic levels.
+One example for the former is experience synthesis: at each RL step, the agent (backed by the rollout LLM) iteratively generates refined responses to a query by incorporating feedback or guidance from the environment, which can be in the form of plain text rather than numerical reward values.
+The resulting data will then be utilized for updating the policy model, e.g., by a standard SFT or RL loss.
+Such a macroscopic RL approach is made possible by pre-trained LLMs’ generative nature and rich prior knowledge about natural language.
+Closely related to this idea is Agent-RLVR [4], a contemporary work that applies such an approach to software engineering scenarios.
+Within Trinity-RFT, this process of experience synthesis can be regarded as a particular way of agent-environment interaction, and thus can be realized by simply implementing a Workflow class.
+As a minimal demonstration, suppose that we want to implement this approach for a math reasoning scenario, where the agent generates multiple rollout responses to an input query, receives feedback from the environment regarding correctness of the responses, reflects on the gathered information, and generates a final response to the query.
+Listing 3 presents an implementation of this approach within Trinity-RFT.
+1@WORKFLOWS.register_module("reflect_once_workflow")
+2class ReflectOnceWorkflow(Workflow):
+3
+4 def run(self) -> List[Experience]:
+5 experiences = []
+6
+7 # Stage 1: K-rollout generation
+8 rollout_messages = self.create_rollout_messages()
+9 responses = self.model.chat(
+10 rollout_messages,
+11 n=self.k_rollouts,
+12 temperature=self.temperature,
+13 logprobs=self.logprobs,
+14 max_tokens= self.task.rollout_args.max_tokens,
+15 )
+16 rollout_responses = [response.response_text.strip() for response in responses]
+17
+18 # Stage 2: Verification
+19 verification_results = []
+20 for rollout_response in rollout_responses:
+21 is_correct = self.verify_answer(rollout_response, self.ground_truth)
+22 verification_results.append(is_correct)
+23
+24 # Stage 3: Reflection
+25 reflection_messages = self.create_reflection_messages(
+26 rollout_responses,
+27 verification_results,
+28 )
+29 reflection_responses = self.model.chat(
+30 reflection_messages,
+31 n=1,
+32 temperature=self.temperature,
+33 logprobs=self.logprobs,
+34 max_tokens=self.task.rollout_args.max_tokens,
+35 )
+36 reflection_response = reflection_responses[0]
+37
+38 # Verify the reflection response
+39 reflection_text = reflection_response.response_text.strip()
+40 reflection_is_correct = self.verify_answer(reflection_text, self.ground_truth)
+41
+42 if reflection_is_correct:
+43 sharegpt_message = [
+44 {
+45 "role": "system",
+46 "content": self.task.format_args.system_prompt
+47 },
+48 {
+49 "role": "user",
+50 "content": self.question
+51 },
+52 {
+53 "role": "assistant",
+54 "content": reflection_text
+55 }
+56 ]
+57 experience = self.process_messages_to_experience(sharegpt_message)
+58 experiences.append(experience)
+59
+60 # Save experience to file
+61 if self.exp_file and sharegpt_message is not None:
+62 exp_data = sharegpt_message
+[…]
+64 self.exp_file.flush()
+65 return experiences
+Listing 3: A toy implementation of experience synthesis with environmental feedback.
+3.2 RL Algorithm Development with Trinity-RFT
+To support RL algorithm development, Trinity-RFT allows researchers and developers to focus on designing and implementing the essential logic of a new RL algorithm, without the need to care about the internal engineering details about Trinity-RFT.
+As an example, suppose that we want to implement a MIX algorithm that seamlessly integrates online RL and offline SFT into a single learning process.
+In its most basic form, the MIX algorithm requires that (1) the trainer samples from two sources of experiences, i.e., the rollout experiences collected online and the high-quality expert trajectories collected offline; and (2) the trainer updates its policy model with a loss function that handles both sources of experiences properly, e.g., a weighted sum of GRPO loss for the on-policy rollout experiences and SFT loss for the expert trajectories.
+Variants of this MIX algorithm include adaptive weighting of multiple loss terms [9], alternating between RL and SFT [15], incorporating expert trajectories into RL loss [20, 33, 45], or incorporating SFT loss for high-reward rollout trajectories generated by older versions of the rollout model [26].
+Such approaches have proved to be effective in accelerating the online RL process with only a small amount of expert experiences, or to enhance stability and plasticity in continual learning.
+Figure 8: A visualization of the MIX algorithm.
+The MIX algorithm is visualized in Figure 8, where we integrate GRPO loss for usual experiences generated by the rollout model and SFT loss for expert experiences into a unified training pipeline.
+It requires dealing with different sources of experiences, and two types of loss functions; fortunately, to implement such an algorithm in Trinity-RFT, we only need to define three new classes — MixSampleStrategy, MIXPolicyLossFn, and MIXAlgorithm — as demonstrated in Listing 4.
+With these components, Trinity-RFT enables a seamless integration of online RL and offline SFT within the same training loop.
+More details of the MIX algorithm are referred to the documentation [^5] https://modelscope.github.io/Trinity-RFT/tutorial/example_mix_algo.html.
+1@SAMPLE_STRATEGY.register_module("mix")
+2class MixSampleStrategy(SampleStrategy):
+3 def __init__(self, buffer_config: BufferConfig, trainer_type: str, **kwargs):
+4 # ...
+5 self.usual_exp_buffer = get_buffer_reader(
+6 buffer_config.trainer_input.experience_buffer, usual_buffer_config
+7 )
+8 self.expert_exp_buffer = get_buffer_reader(
+9 buffer_config.trainer_input.sft_warmup_dataset, expert_buffer_config
+10 )
+11 # ...
+12
+13 def sample(self, step: int) -> Tuple[Any, Dict, List]:
+14 """Sample a batch composed of rollout experiences and expert trajectories"""
+15 usual_exp_list = self.usual_exp_buffer.read()
+16 expert_exp_list = self.expert_exp_buffer.read()
+17 exp_list = usual_exp_list + expert_exp_list
+18 exps = Experiences.gather_experiences(exp_list, self.pad_token_id)
+19 # ...
+20
+21
+22@POLICY_LOSS_FN.register_module("mix")
+23class MIXPolicyLossFn(PolicyLossFn):
+24 def __init__(self, mu: float = 0.1, ...):
+25 # ...
+26 self.mu = mu
+27 self.grpo_loss_fn = PPOPolicyLossFn(...)
+28 self.sft_loss_fn = SFTLossFn(...)
+29
+30 def __call__(
+31 self,
+32 logprob: torch.Tensor,
+33 old_logprob: torch.Tensor,
+34 action_mask: torch.Tensor,
+35 advantages: torch.Tensor,
+36 is_expert_mask: torch.Tensor,
+37 **kwargs,
+38 ) -> Tuple[torch.Tensor, Dict]:
+39 """Calculate a weighted sum of GRPO loss and SFT loss"""
+40 # ...
+41 grpo_loss, grpo_metrics = self.grpo_loss_fn(
+42 logprob[~is_expert_mask],
+43 old_logprob[~is_expert_mask],
+44 action_mask[~is_expert_mask],
+45 advantages[~is_expert_mask],
+46 **kwargs,
+47 )
+48 sft_loss, sft_metrics = self.sft_loss_fn(
+49 logprob[is_expert_mask],
+50 action_mask[is_expert_mask],
+51 )
+52 loss = (1 - self.mu) * grpo_loss + self.mu * sft_loss
+53 # ...
+54 return loss, metrics
+55
+56@ALGORITHM_TYPE.register_module("mix")
+57class MIXAlgorithm(AlgorithmType):
+58 """MIX algorithm."""
+59
+60 use_critic: bool = False
+61 use_reference: bool = True
+62 use_advantage: bool = True
+63 can_balance_batch: bool = True
+64 schema: type = ExperienceModel
+65
+66 @classmethod
+67 def default_config(cls) -> Dict:
+68 return {
+69 "repeat_times": 8,
+70 "policy_loss_fn": "mix", # Specify the MIX loss function
+71 "advantage_fn": "grpo",
+72 "sample_strategy": "mix", # Specify the MIX sampling strategy
+73 }
+Listing 4: An implementation of the MIX algorithm with Trinity-RFT.
+3.3 Unified Support for Diverse RL Modes
+As explained previously in Section 2.1.1, Trinity-RFT offers support for synchronous/asynchronous, on-policy/off-policy, and online/offline RL, controlled by a few configuration parameters.
+In this subsection, we conduct experiments for comparing the following RL modes:
+  • The synchronous mode: mode=both, sync_interval={1,2,10}, sync_offset=0;
+  • The one-step off-policy mode: mode=both, sync_interval=1, sync_offset=1;
+  • The fully asynchronous mode: the explorer and trainer are launched with mode=explore and train respectively, with sync_interval=10.
+Our experiments include dummy learning processes (which will soon be explained) for performance profiling, as well as real learning processes with vanilla GRPO [28] in different modes.
+3.3.1 Experiments: Performance Profiling
+Settings.
+We aim to measure and compare the efficiency of different RL modes under controlled settings.
+It is noteworthy that, even with all other variables controlled, different RL modes can still result in different trained models — and thus different rollout response lengths — during the learning processes, which have direct impacts on performance metrics like wall-clock time and GPU utilization rate.
+To mitigate this, we conduct performance profiling with dummy learning processes, where the learning rate is set to zero.
+A dummy learning process closely resembles a real learning process, in that all necessary computation and communication (e.g., rollout generation, gradient computation, model weight synchronization) are executed; the only difference is that the rollout model (and thus the distribution of rollout trajectories) remains fixed throughout and same across different RL modes.
+To show the performance of Trinity-RFT in diverse scenarios, we consider both math reasoning task (GSM8k [3]) and multi-turn agentic task (ALFWorld [30]).
+In the experiments, we use the Qwen2.5-Instruct [35] model series of different sizes (1.5B, 3B, and 7B), and run the GRPO [5] algorithm (with 8 rollout trajectories per task) in all modes.
+We choose a 100-step training trace to evaluate the performance and report the following metrics: (1) end-to-end wall-clock time and time speedup: the wall-clock time from the start of running the command to the end of finishing 100 training steps; (2) GPU utilization: the GPU utilization in percent for each GPU; (3) GPU power usage: the GPU power usage as a percentage of its power capacity for each GPU.
+Metrics for GPU utilization and power usage were extracted from WandB [^6] https://docs.wandb.ai/guides/models/app/settings-page/system-metrics/ and averaged over all GPUs.
+We run each experiment for three random trials and report the mean and standard deviation.
+Unless specified otherwise, each experiment uses 8 NVIDIA A100-80G GPUs.
+Results for GSM8k.
+We use the 2/6 GPU partition for explorer/trainer.
+While this configuration is not the optimal one for all experiments, it is sufficient to show the difference between several RL modes.
+In our GSM8k experiments, the batch size is 96 tasks, and the temperature is 1.0.
+The results for both Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct models are shown in Table 1.
+We observe that in the synchronous mode (with sync_offset=0), a less frequent synchronization (i.e., a larger sync_interval) effectively improves the efficiency, GPU utilization, and GPU power usage.
+This is mainly because, as shown in Figure 4 (a), the impacts of pipeline bubbles in this mode can be effectively reduced by using a lower synchronization frequency.
+Besides, Table 1 shows that one-step off-policy and fully asynchronous modes also accelerate the training process with higher GPU utilization, compared to a strictly on-policy mode.
+In one-step off-policy mode, the trainer leverages the one-step off-policy data stored in the buffer without needing to wait for new experiences generated by the explorer after weight synchronization, which significantly reduces the GPU idle ratio.
+In fully asynchronous mode, the explorer and trainer operate almost independently while fully leveraging GPU resources, except when loading or saving model checkpoints.
+Table 1: Performance profiling for GSM8k with 2/6 GPU partition for explorer/trainer.
+Mode Qwen2.5-1.5B-Instruct
+Speedup $\uparrow$ Time (Minutes) $\downarrow$ GPU Utilization (%) $\uparrow$ GPU Power Usage (%) $\uparrow$
+Sync. (sync_interval=1) 1.00×1.$0$ $38.70\pm 0.34$ $33.64\pm 2.15$ $35.85\pm 1.83$
+Sync. (sync_interval=2) 1.24×1.24$\times$1.24 × $31.19\pm 0.08$ $36.05\pm 0.49$ $38.74\pm 1.47$
+Sync. (sync_interval=10) 1.59×1.59$\times$1.59 × $24.28\pm 0.16$ $\mathbf{38.27}\pm 0.98$ $\mathbf{44.41}\pm 0.81$
+One-step off-policy 1.25×1.25$\times$1.25 × $30.84\pm 0.20$ $32.39\pm 1.17$ $39.70\pm 0.78$
+Fully async. $\mathbf{1.61}\times$ $\mathbf{23.97}\pm 0.03$ $36.04\pm 0.61$ $43.91\pm 0.48$
+Mode Qwen2.5-7B-Instruct
+Speedup $\uparrow$ Time (Minutes) $\downarrow$ GPU Utilization (%) $\uparrow$ GPU Power Usage (%) $\uparrow$
+Sync. (sync_interval=1) 1.00×1.$0$ $68.71\pm 0.54$ $55.61\pm 0.80$ $52.88\pm 0.29$
+Sync. (sync_interval=2) 1.31×1.31$\times$1.31 × $52.44\pm 0.41$ $64.88\pm 1.35$ $61.90\pm 1.32$
+Sync. (sync_interval=10) $\mathbf{1.85}\times$ $\mathbf{37.17}\pm 0.15$ $\mathbf{78.44}\pm 1.03$ $\mathbf{77.77}\pm 0.96$
+One-step off-policy 1.69×1.69$\times$1.69 × $40.73\pm 0.57$ $77.19\pm 2.26$ $76.17\pm 1.56$
+Fully async. 1.63×1.63$\times$1.63 × $42.17\pm 1.06$ $73.90\pm 2.00$ $72.74\pm 1.82$
+Table 2: Performance profiling for ALFWorld with 4/4 GPU partition for explorer/trainer.
+Mode Batch_Size = 4
+Speedup $\uparrow$ Time (Minutes) $\downarrow$ GPU Utilization (%) $\uparrow$ GPU Power Usage (%) $\uparrow$
+Sync. (sync_interval=1) 1.00×1.$0$ $333.68\pm 0.06$ $17.19\pm 0.58$ $28.44\pm 0.37$
+Sync. (sync_interval=2) 1.70×1.70$\times$1.70 × $196.64\pm 0.59$ $21.69\pm 0.18$ $31.35\pm 0.06$
+Sync. (sync_interval=10) 5.21×5.21$\times$5.21 × $64.09\pm 0.39$ $32.85\pm 0.18$ $40.86\pm 0.58$
+One-step off-policy 0.98×0.98$\times$0.98 × $340.12\pm 3.99$ $14.63\pm 1.17$ $28.21\pm 0.48$
+Fully async. $\mathbf{5.45}\times$ $\mathbf{61.27}\pm 0.35$ $\mathbf{36.46}\pm 0.10$ $\textbf{42.51}\pm 0.72$
+Mode Batch_Size = 32
+Speedup $\uparrow$ Time (Minutes) $\downarrow$ GPU Utilization (%) $\uparrow$ GPU Power Usage (%) $\uparrow$
+Sync. (sync_interval=1) 1.00×1.$0$ $561.21\pm 2.04$ $39.37\pm 0.89$ $39.93\pm 0.22$
+Sync. (sync_interval=2) 1.13×1.13$\times$1.13 × $496.80\pm 0.36$ 37.74 ±$\pm 0.39$ $41.90\pm 0.44$
+Sync. (sync_interval=10) 1.59×1.59$\times$1.59 × $352.11\pm 0.49$ $44.50\pm 0.58$ $49.95\pm 0.61$
+One-step off-policy 1.14×1.14$\times$1.14 × $494.13\pm 0.28$ $34.89\pm 0.75$ $43.05\pm 0.81$
+Fully async. $\mathbf{1.65}\times$ $\mathbf{339.51}\pm 0.24$ $\mathbf{45.55}\pm 0.20$ $\textbf{50.77}\pm 0.45$
+Results for ALFWorld.
+A particular feature of ALFWorld is the long-horizon multi-turn interaction with the environment.
+To accommodate the heavy computational demands in rollout, we use the 4/4 GPU partition for explorer/trainer.
+In our ALFWorld experiments, we use the Qwen2.5-3B-Instruct model and set the rollout temperature to 1.0.
+The results with different batch sizes are shown in Table 2.
+One observation is that, when the batch size is 4 tasks, the one-step off-policy mode exhibits no efficiency advantage over the synchronous mode with sync_interval=1.
+This phenomenon can be attributed to the computational imbalance between the explorer and trainer.
+In ALFWorld, the larger computation latency of the explorer emerges primarily from (1) the inherent complexity of multi-turn environment interactions, and (2) the long-tailed latency distribution when certain tasks require extended rollout durations, whose effect is further exacerbated by a small batch size.
+The one-step off-policy mode cannot eliminate pipeline bubbles caused by long-tailed latencies in the explorer, whereas this can be mitigated by the synchronous mode with a large sync_interval as well as by the asynchronous mode, thanks to the implementation of streaming rollout generation in Trinity-RFT.
+Another observation, due to the same reason, is that when scaling the batch size from 4 to 32, all modes incur a small increase (much smaller than 8×8$\times$8 ×) in wall-clock time for the same number of training steps, thanks to better GPU usage.
+3.3.2 Experiments: Real Learning with Vanilla GRPO
+Settings.
+We aim to compare the real learning processes by different RL modes.
+For simplicity and controlled variability, we use the vanilla GRPO [28] algorithm for all RL modes, without specific algorithm design for asynchronous or off-policy cases.
+GRPO mainly relies on the mechanism of clipping probability ratio (defaults to the range of $1\pm 0.2$) to handle the off-policyness of experiences.
+For future works, we will investigate more advanced off-policy or asynchronous RL algorithms (e.g., the OPMD algorithm proposed by Kimi k1.5 [34] and two OPMD variants developed in Appendix A), and develop new ones to accommodate diverse RL modes.
+To encourage the exploration of the rollout model, we disable the Kullback-Leibler (KL) penalty or loss in our experiments.
+Training.
+For each RL mode, we fine-tune the Qwen2.5-7B-Instruct model for one epoch on the OpenR1-Math-46k [46] [^7] https://huggingface.co/datasets/Elliott/Openr1-Math-46k-8192 dataset, a filtered version of the OpenR1-Math-220k [^8] https://huggingface.co/datasets/open-r1/OpenR1-Math-220k dataset.
+The allocated GPU ratio for the explorer and trainer is 4/4.
+We set the batch size to 120 tasks, the rollout number per task to 32, and the learning rate to 1e-6.
+Evaluation.
+For each RL mode, we save the checkpoint of the rollout model once every 100 steps, evaluate the checkpoints using the bench mode, and report the best results among the checkpoints.
+The models are evaluated on several math benchmarks, including AIME2024999https://huggingface.co/datasets/math-ai/aime25, AIME2025101010https://huggingface.co/datasets/math-ai/aime25, AMC [^11] https://huggingface.co/datasets/math-ai/amc23, and MATH500121212https://huggingface.co/datasets/HuggingFaceH4/MATH-500.
+For AIME2024, AIME2025, and AMC, we generate 32 responses (with temperature 0.6) per task and report the average accuracy (Avg@32), to ensure the accuracy of the evaluation process; for MATH500, we report Avg@4 as the dataset is relatively large.
+For more detailed comparisons, we also plot some training metrics, including reward, response length, gradient norm, and KL distance from the initial LLM, with the wall-clock time as the X-axis.
+Results.
+Figure 9 presents the training curves.
+It is observed that several RL modes show increasing trends in terms of rewards and response lengths.
+The synchronous mode with sync_interval=1 exhibits longer responses and larger KL divergence than other RL modes, likely because it updates the rollout model most frequently and leverages on-policy data in each step.
+Table 3 presents the evaluation results.
+We observe that, for the synchronous mode with sync_offset=0, increasing sync_interval reduces the total training time for one epoch, at the cost of slightly compromising the average evaluation performance.
+In contrast, the one-step off-policy mode with sync_interval=1 achieves comparable or better performance than the other modes on several benchmarks, while achieving around 2.66×2.$6$ speedup in wall-clock time over the strictly on-policy mode.
+Figure 9: Results of training for one epoch by vanilla GRPO in different RL modes. The results are smoothed using a 40-step moving average for clarity.
+Table 3: Performance comparison of different RL modes.
+AIME2024 AIME2025 AMC MATH500 Average Runtime (Hours)
+Qwen2.5-7B-Instruct 11.15 6.95 51.35 70.96 35.10 N/A
+Sync. (sync_interval=1) 14.58 14.06 61.25 76.25 41.54 130.33
+Sync. (sync_interval=2) 15.52 11.67 57.97 75.15 40.08 73.57
+Sync. (sync_interval=10) 14.38 12.71 57.66 75.05 39.95 44.43
+One-Step Off-Policy 16.88 12.19 59.92 74.55 40.89 48.98
+3.4 Data Processors for Tasks and Experiences
+We present practical examples to demonstrate how the data pipeline concepts described in Section 2.3 are applied in Trinity-RFT.
+These use cases highlight how manipulating data at both the task and experience level directly improves RFT performance and provides granular control over the agent’s learning process.
+3.4.1 Static Task Prioritization for Curriculum Learning
+A common strategy for effective training is to present tasks in increasing order of difficulty.
+This use case demonstrates how Trinity-RFT facilitates curriculum learning by prioritizing tasks before exploration begins.
+This is particularly crucial for RFT, as it helps stabilize the initial learning phase of the explorer and prevents it from getting stuck on overly complex tasks, leading to a more efficient exploration path.
+As shown in Listing 5, a user can configure this pipeline with a simple YAML file  [^13] The full configuration files can be accessed at baseline_run and priority_run..
+Here, we use the GSM8K mathematical reasoning dataset.
+The user provides a natural language instruction via dj_process_desc: “Please compute difficulty scores for these math questions.”.
+Trinity-RFT’s data service then orchestrates a three-phase process:
+  1. 1.
+    The data processor invokes an LLM (Qwen-Max) to score the difficulty of each math problem.
+  2. 2.
+    The system prioritizes samples with lower difficulty scores, creating an easy-to-hard ordering (by setting ‘priority_weights["difficulty"]: -1.0’).
+  3. 3.
+    The curated and prioritized data is formatted into an RL-ready task set for the explorer.
+As shown in Figure 10, this simple curation strategy yields more stable performance gains.
+This pattern is highly extensible: users can easily customize the difficulty metric, apply it to their own datasets, or even make the prioritization dynamic by re-ranking tasks periodically based on the agent’s current performance.
 1# Core dataset configuration
-2class DataConfig:
-3 """Centralized dataset configuration management"""
-4
-5 def __init__(self):
-6 self.data = {
-7 "dataset_path": "/path/to/gsm8k",
-8 "format_config": {
-9 "prompt_key": "question",
-10 "response_key": "answer"
-11 },
-12 "dj_process_desc": "Compute difficulty scores for math questions",
-13 "agent_model_name": "qwen-max"
-14 }
-15
-16 def get_metadata(self) -> dict:
-17 """Retrieve configuration metadata"""
-18 return self.data
-Listing 3: Core dataset configuration, which includes processing instructions as metadata.
-1# Typical procedure that transforms raw dataset through multi-stage pipeline
-2my_dataset = RftDataset(data_config) # Raw dataset auto-loading
-3
-4my_dataset.format([
-5 BoxedMathAnswerFormatter(config), # Domain-specific formatting
-6 RLHFFormatter(config) # RLHF format conversion
-7 ])
-8
-9my_dataset.to_taskset(
-10 reward_fn=AccuracyReward, # Reward shaping specification
-11 workflow=MathWorkflow # Execution environment binding
-12 )
-Listing 4: Dataset transformation with domain-specific formatting and RLHF conversion.
-4.4 Human-AI Collaborative Annotation
+2data_processor:
+3  data_workflow_url"http://127.0.0.1:5005/data_workflow"
+4 task_pipeline:
+5  # I/O buffers
+6 input_buffers:
+7  - name"raw_input"
+8 path"openai/gsm8k"
+9 storage_type"file"
+10 rawtrue
+11 output_buffer:
+12  name"raw_output"
+13 path"outputs/task_pipeline_output/prioritized_gsm8k.jsonl"
+14 storage_type"file"
+15 format:
+16  prompt_key"question"
+17 response_key"answer"
+18 # data active iterator related
+19 dj_process_desc"Please compute difficulty scores for these math questions."
+20 agent_model_name"qwen-max"
+21 clean_strategy"iterative"
+22 priority_weights:
+23  difficulty-1.0 # easy-to-hard
+Listing 5: Data processor configuration, applied on customizable buffers.
+Figure 10: Performance on a math reasoning task. Prioritizing tasks from easy to hard (red line) leads to faster and better convergence compared to the default setting (blue line).
+3.4.2 Dynamic Experience and Reward Shaping
+While task curation primes the model before exploration, experience shaping refines the learning signal after each agent-environment interaction.
+This is critical for RFT algorithms that rely on rich feedback, as standard rewards (e.g., binary pass/fail) are often too sparse to guide learning effectively.
+We demonstrate how to augment rewards with metrics for quality and diversity, transforming a sparse signal into a dense, multi-faceted one that provides clearer guidance to the trainer.
+Use Case 1: Quality Reward Augmentation.
+To encourage the model to generate high-quality responses, we can augment the base reward with a quality score.
+As illustrated in Figure 11, during each RFT step, we use the data processor to evaluate the quality of each generated rollout.
+For our experiment, we trained a Qwen2.5-1.5B model and used a more powerful Qwen3-32B as the scorer LLM.
+Specifically, we invoked the llm_quality_filter from Data-Juicer, which normalized the quality scores to the range [-0.5, 0.5] and added them to the original reward.
+Crucially, this processing is applied to the experience buffer at each RFT step.
+This allows the reward signal to adapt dynamically to the policy model’s evolving capabilities, a more responsive approach than one-time static processing.
+With a sync_interval of 3 over 36 steps on the Math-500 validation set, the results in Figure 12 show that: (1) The model with quality reward augmentation (red line) achieves higher accuracy.
+(2) The introduced quality reward itself improves over time, confirming it is a learnable signal.
+(3) We observe a slight increase in response length, which likely reflects an inductive bias from the larger scorer model being implicitly distilled into the smaller policy model.
+Figure 11: The enhanced math workflow with quality-reward shaping from data processor, where DJ indicates DataJuicer [2], from which more operators can be utilized to extend this worklow.
+Figure 12: Experimental results for quality-reward shaping. Augmenting the reward with a quality score (red line) improves final accuracy and provides a learnable reward signal.
+Use Case 2: Diversity Reward Augmentation.
+A common failure mode in RFT is policy collapse, where the agent repeatedly generates similar, suboptimal responses.
+To counteract this, we introduce a diversity reward that encourages the explorer to explore different solution paths.
+As shown in Figure 13, we used the GTE-Qwen2-1.5B model to convert rollouts into semantic embeddings.
+The diversity reward was calculated based on the cosine similarity of a rollout’s embedding to the mean embedding of its group, with lower similarity (i.e., higher diversity) yielding a higher reward.
+To prevent exploration from becoming chaotic, we applied a simple decay schedule to the diversity reward’s weight, starting at 0.5 and decaying to 0.3 over the training steps.
+The experiment, using the same setting as before, yielded compelling results (Figure 14): (1) The diversity-augmented model (red line) shows a significant performance improvement over the baseline.
+(2) The response length is consistently longer, indicating the reward encourages more elaborate answers.
+(3) Most importantly, the actor entropy loss remains consistently higher, providing strong evidence that the model is maintaining a healthier, more diverse exploration strategy, which helps it escape local optima.
+Figure 13: The enhanced math workflow with diversity-reward shaping from data processor
+Figure 14: Experimental results for diversity-reward shaping. Rewarding diverse responses (red line) significantly improves task accuracy and maintains higher entropy.
+3.5 RFT with Human in the Loop
 This example demonstrates the human-in-the-loop capability in Trinity-RFT for preference modeling.
-As illustrated in Listing 5 and Figure 5, our framework integrates Label Studio’s annotation interface with asynchronous data pipelines through four coordinated stages: (1) task generation: auto-creating annotation batches from model rollouts; (2) interactive labeling: providing UI for side-by-side response comparison; (3) quality control: enforcing inter-annotator agreement thresholds; and (4) versioned storage: tracking preference lineage in pre-defined fields like those in DPODataModel.
+As illustrated in Listing 6 and Figure 15, our framework integrates Label Studio’s annotation interface with asynchronous data pipelines through four coordinated stages: (1) task generation: auto-creating annotation batches from model rollouts; (2) interactive labeling: providing UI for side-by-side response comparison; (3) quality control: enforcing inter-annotator agreement thresholds; and (4) versioned storage: tracking preference lineage in pre-defined fields like those in DPODataModel.
 This pipeline reflects Trinity-RFT’s bi-directional collaboration feature (Section 2.3.4), backed by timeout-aware task polling and support of atomic batch commit.
@@
 20 return self.process
-Listing 5: Configuration for human preference annotation.
+Listing 6: Configuration for human preference annotation.
 Figure 15: An interactive interface for human preference annotation.
-4.5 Trinity-Studio
+3.6 Low-Code Usage and Development with Trinity-Studio
+(a) Trinity-Studio dashboard.
+(b) Start training on the “Training Portal” page.
+(c) Manage data on the “pgAdmin” page.
+(d) Process data on the “Label Studio” page.
+Figure 16: Snapshots of Trinity-Studio.
 Trinity-Studio provides visual interaction for the core capabilities of Trinity-RFT, designed to bridge the gap between system complexity and user accessibility.
-As shown in Figure LABEL:fig:trinity-studio:dashboard, its three integrated modules, “Training Portal”, pgAdmin”, and “Label Studio”, form a cohesive interface that directly supports the active data-reward co-shaping and human-AI collaboration capabilities established in Section 3.3.
-  • “Training Portal” (Figure LABEL:fig:trinity-studio:training) implements configuration-to-execution procedures through declarative YAML editing with live validation and auto completion.
-This design aligns with Trinity-RFT’s decoupled architecture, enabling users to optionally specify dataset paths, processing and training parameters, meanwhile ensuring schema-aware validation against pre-defined data schema constraints (Section B.1).
-The live validation mechanism prevents misconfigurations that could disrupt the explorer-trainer coordination described in Section 3.1.
+As shown in Figure 16(a), its three integrated modules — “Training Portal”, “pgAdmin”, and “Label Studio” — form a cohesive interface that supports low-code usage and development with Trinity-RFT, and make it easy to monitor and track the full RFT pipeline with transparency.
+  • “Training Portal” (Figure 16(b)) implements configuration-to-execution procedures through declarative YAML editing with auto completion and live validation that prevents misconfigurations.
 Furthermore, the integration of runtime metrics with tools like Wandb/TensorBoard directly helps the active data optimization feature by surfacing signals such as difficulty distribution drifts and diversity metrics mentioned in Section 3.4.
 This transparency ensures that users can monitor how data curation strategies impact RFT performance in real time.
-  • “pgAdmin” (Figure LABEL:fig:trinity-studio:pgadmin) reflects Trinity-RFT’s end-to-end data transformation capabilities by providing a visual panel for PostgreSQL-based storage (Section 3.3.1).
+  • “pgAdmin” (Figure 16(c)) reflects Trinity-RFT’s end-to-end data transformation capabilities by providing a visual panel for PostgreSQL-based storage.
 This design benefits the versioned data lineage requirements of RFT, particularly for scenarios involving asynchronous training (Section 2.3.3).
@@
 This capability is valuable for rapid validation of active learning policies by cross-referencing training outcomes with metadata (e.g., difficulty scores and staleness in asynchronous mode).
-  • “Label Studio” page (Figure LABEL:fig:trinity-studio:label-studio) operationalizes Trinity-RFT’s bi-directional human-AI collaboration capability (Section 3.3.3).
+  • “Label Studio” page (Figure 16(d)) operationalizes Trinity-RFT’s bi-directional human-AI collaboration capability (Section 2.3.4).
 Utilizing the provided task polling and atomic batch commit mechanisms, users can annotate the data or experiences directly, allowing an asynchronous way to involve human feedback and to dynamically influence data curation.
@@
 For example, a researcher tuning a math reasoning task could use the Training Portal to adjust difficulty scoring parameters, view the resulting distribution shifts in the pgAdmin module, and then validate human annotators’ preferences in the Label Studio page.
-This end-to-end visibility is critical for debugging and iterating on RFT strategies, particularly when balancing tradeoffs between quality, diversity, and reward shaping.
-Moreover, this visual operation complements the programmatic APIs of Trinity-RFT while maintaining full compatibility with CLI procedures.
-We implement Trinity-Studio with Singe-Spa framework [27].
-The modular architecture enables custom view development through JavaScript plugins and flexible extensions for general-purpose usage.
-Figure 6: Snapshots of Trinity-Studio.
-5 Conclusion
-This report has provided an overview of Trinity-RFT, a general-purpose, flexible, scalable and user-friendly framework for reinforcement fine-tuning of large language models.
-Trinity-RFT offers a path into “the era of experience” [26], by supporting applications in diverse scenarios with complex agent-environment interaction, and serving as a unified platform for exploring advanced RL paradigms.
+This end-to-end visibility can be useful for debugging and iterating RFT strategies, and complements the programmatic APIs of Trinity-RFT while maintaining full compatibility with CLI procedures.
+We implement Trinity-Studio with the Singe-Spa framework [32]. The modular architecture enables custom view development through JavaScript plugins and flexible extensions for general-purpose usage.
+4 Conclusion and Next Steps
+We have presented Trinity-RFT, a general-purpose, flexible, scalable and user-friendly framework for reinforcement fine-tuning of large language models.
+Trinity-RFT offers a path into “the era of experience” [31], by supporting applications in diverse scenarios with complex agent-environment interaction, and serving as a unified platform for exploring advanced methodologies in each stage of the complete RFT pipeline, at both macroscopic and microscopic levels.
+Further development of Trinity-RFT includes enhancing support for generic multi-step agentic workflows, incorporating more advanced RL algorithms (especially off-policy or asynchronous ones), making the choices of hyperparameters more adaptive and less reliant on manual tuning, augmenting data pipelines with smarter sampling strategies and data processing operations, and conducting more thorough experiments and evaluations with Trinity-RFT.
 Acknowledgements
-Trinity-RFT is built upon many excellent open-source projects, including but not limited to: verl [24] and PyTorch’s FSDP [5] for LLM training; vLLM [10] for LLM inference; Data-Juicer [2] for data processing functionalities; AgentScope [6] for agentic workflow; and Ray [14] for distributed systems.
-We have also drawn inspiration from RL/RFT projects like OpenRLHF [8], TRL [31], ChatLearn [1], and TinyZero [19].
+Trinity-RFT is built upon many excellent open-source projects, including but not limited to: verl [29] and PyTorch’s FSDP [7] for LLM training; vLLM [14] for LLM inference; Data-Juicer [2] for data-related functionalities; AgentScope [10] for agentic workflow; and Ray [19] for distributed runtime.
 Appendix A Off-Policy Reinforcement Learning Algorithms
@@
 At the end of this section, we arrive at the surprising conclusion that the standard policy gradient, weighted by a coefficient and using the group mean reward as the baseline, can be a feasible direction for updating the policy even in off-policy settings (while the standard theory of policy gradient only holds for on-policy settings).
-This has been validated empirically in our exploratory experiments during the development of Trinity-RFT.
+This has been validated empirically in our exploratory experiments during the development of Trinity-RFT  [^14] https://github.com/modelscope/Trinity-RFT/blob/main/docs/sphinx_doc/source/tutorial/example_reasoning_advanced.md.
 A.1 OPMD: Kimi’s Version
@@
 A larger $\tau$ is effectively the same as a smaller learning rate.
-As a heuristic, we simply add a regularization term (denoted by $g$) to the above objective when additional regularization with respect to a fixed policy, e.g., a SFT model $\pi_{\textsf{sft}}$, is desired:
+Remark 2.
+As a heuristic, when additional regularization with respect to a fixed policy (e.g., a SFT model $\pi_{\textsf{sft}}$) is desired, we simply add a regularization term (denoted by $g$) to the above objective:
 $\min_{\bm{\theta}}\quad\widehat{J}(\bm{\theta};x)\coloneqq-\frac{1}{1+\tau}\sum_{1\leq i\leq K}\big{(}r(x,y_{i})-\overline{r}(x)\big{)}\log\pi_{\bm{\theta}}(y_{i}|x)+\beta\cdot g\big{(}\pi_{\bm{\theta}},\pi_{\textsf{sft}};x,y_{1},\dots,y_{K}\big{)}.$
-Appendix B Additional Details for Data Pipelines
-B.1 Illustration of Typical Dataset Classes and Schemas
-The dataset shaping process described in Section 3.3 is supported by PostgreSQL persistence with comprehensive schema designs; one example can be found in Listing 6.
-Key model classes include RftDatasetModel (core dataset storage with lineage tracking), ExperienceModel (which stores serialized experiences with priority scoring), DPODataModel and SFTDataModel (specialized storage for DPO and SFT, respectively).
-1-- RftDataset table structure
-2CREATE TABLE rft_dataset (
-3 id SERIAL PRIMARY KEY, -- Unique record identifier
-4 consumed_cnt INTEGER DEFAULT 0, -- Usage counter for sampling
-5 prompt TEXT, -- Input instruction
-6 response TEXT, -- Model-generated response
-7 reward FLOAT, -- Training signal value
-8 quality_score FLOAT DEFAULT 0.0, -- Quality assessment metric
-9 difficulty_score FLOAT DEFAULT 0.0, -- Complexity estimation
-10 diversity_score FLOAT DEFAULT 0.0, -- Response uniqueness measure
-11 priority FLOAT DEFAULT 0.0 -- Sampling preference weight
-12);
-Listing 6: Schema definition for RftDataset PostgreSQL table.
-B.2 Illustration of Data Processing Service
-The architecture presented in Section 3.3 integrates Data-Juicer capabilities through service-oriented design.
-Typically, the DJ-service lifecycle covers three stages: initialization via scripts/start_servers.py; configuration through dj_config_path or natural language (dj_process_desc); and execution handling dataset loading, cleaning, priority scoring, and database export.
-Moreover, we provide DJExecutor to bridge Data-Juicer capabilities through config translation and execution sandboxing, and dynamic termination via _check_clean_targets() with melt protection.
-B.3 An Example of Active Data Optimization Algorithms
-The DataActiveIterator serves as the core engine for data-model co-evolution with multi-dimensional utility scoring.
-It provides extensible interfaces for flexible training prior injection, e.g., as demonstrated in Algorithm 1.
-Algorithm 1 Multi-Dimensional Sample Selection
-0:  Dataset $D$, weights $w_{q},w_{d},w_{k},w_{f}$
-0:  Selected subset $D^{\prime\prime}$
-1:  $(q,d,k)\leftarrow\text{DataCleaner}.\text{Process}(D)$ Compute quality ($q$), diversity ($d$), difficulty ($k$)
-2:  $U_{i}\leftarrow w_{q}q_{i}+w_{d}d_{i}+w_{k}k_{i}-w_{f}f_{i}$ Combine $q$, $d$, $k$, frequency ($f$)
-3:  $D^{\prime}\leftarrow\{s_{i}\in D\mid U_{i}>\theta_{\text{min}}\}$ Threshold filtering
-4:  $D^{\prime\prime}\leftarrow\text{Top-K}(D^{\prime},K)$ Priority selection
-5:  $\text{ExperienceCache}.\text{Update}(D^{\prime\prime})$ Update version-controlled cache
-6:  $\text{LineageTracker}.\text{Record}(D^{\prime\prime})$ Track data provenance
-7:  return $D^{\prime\prime}$